Linux NET-2/NET-3 HOWTO

Terry Dawson, terryd@extro.ucc.su.oz.au v2.8, 07 Jan 1995

This document aims to describe how to obtain, install and configure the Linux NET-2 and NET-3 networking software. Some answers to some of the more frequently asked questions are included in the appendix. This document is not designed to teach you about tcp/ip networking, though some information of this kind is included where possible. Pointers to other documentation which does teach tcp/ip networking principles is listed.

Contents

1 Introduction. 5 1.1 Changes from the previous release. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.2 A brief development history of Linux Networking. . . . . . . . . . . . . . . . . . . . . . . . . . 5

2 Disclaimer. 6

3 Questions already ? 7

4 Related Documentation. (Where to learn about tcp/ip) 7 4.1 New versions of this document. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.2 Feedback. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

5 Some terms used in this document. 10

6 NET-2/NET-3 Supported functionality. 11 6.1 Supported Ethernet cards. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

7 Getting the NET-2/NET-3 software. 12 7.1 The kernel source. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7.2 The libraries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7.3 The network configuration tool suite. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 7.4 The network applications. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 7.5 Additional drivers or packages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

8 Configuring the kernel. 16 8.1 What do all those funky Networking options actually do? . . . . . . . . . . . . . . . . . . . . 18 9 Configuring the Network Devices. 20

9.1 Configuring the special device files in /dev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 9.2 What information do I need before I begin ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

9.2.1 IP Address. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 9.2.2 Network Mask (‘netmask’). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 9.2.3 Network Address. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 9.2.4 Broadcast Address. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 9.2.5 Router (‘Gateway’) Address. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 9.2.6 Nameserver Address. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 9.2.7 NOTE for SLIP/PLIP/PPP users. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

9.3 /etc/rc.d/rc.inet1,2 or /etc/rc.net . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 9.3.1 rc.inet1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 9.3.2 rc.inet2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

9.4 Configuring the Loopback device (mandatory). . . . . . . . . . . . . . . . . . . . . . . . . . . 24 9.5 Configuring an ethernet device. (optional) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 9.6 Configuring a SLIP device (optional) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26

9.6.1 dip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 9.6.2 slattach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 9.6.3 When do I use which ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 9.6.4 Static slip server with a dialup line and DIP. . . . . . . . . . . . . . . . . . . . . . . . 28 9.6.5 Dynamic slip server with a dialup line and DIP. . . . . . . . . . . . . . . . . . . . . . . 28 9.6.6 Using DIP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 9.6.7 Permament slip connection using a leased line and slattach. . . . . . . . . . . . . . . . 32

9.7 Configuring a PLIP device. (optional) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 9.7.1 PLIP cabling diagram. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

10 Routing. (mandatory) 34 10.1 Static/Manual Routes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 10.2 Default Route. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 10.3 Proxy ARP. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 10.4 gated - the routing daemon. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

10.4.1 Obtaining gated . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 10.4.2 Installing gated . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 11 Configuring the network daemons. 39

11.1 /etc/rc.d/rc.inet2 (the second half of rc.net) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 11.1.1 inetd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 11.1.2 syslogd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

11.2 A sample rc.inet2 file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 11.3 Other necessary network configuration files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

11.3.1 A sample /etc/inetd.conf file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 11.3.2 A sample /etc/services file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 11.3.3 A sample /etc/protocols file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

11.4 Name Resolution. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 11.4.1 /etc/hosts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 11.4.2 named - do I need thee ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 11.4.3 /etc/networks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 11.4.4 /etc/host.conf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 11.4.5 /etc/resolv.conf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 11.4.6 Configuring your Hostname - /etc/HOSTNAME . . . . . . . . . . . . . . . . . . . . . 49

11.5 Other files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

12 Advanced Configurations. 50

12.1 PPP - Point to Point Protocol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 12.1.1 Why would I use PPP in place of SLIP ? . . . . . . . . . . . . . . . . . . . . . . . . . 51 12.1.2 Where to obtain the PPP software. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 12.1.3 Installing the PPP software. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 12.1.4 Configuring and using the PPP software. . . . . . . . . . . . . . . . . . . . . . . . . . 53 12.1.5 Where to obtain more information on PPP, or report bugs. . . . . . . . . . . . . . . . 55

12.2 Configuring Linux as a Slip Server. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 12.2.1 Slip Server using sliplogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 12.2.2 Slip Server using dip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 12.2.3 slip servers using the dslip package. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

12.3 Using the Automounter Daemon - AMD. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 12.3.1 What is an automounter, and why would I use one ? . . . . . . . . . . . . . . . . . . . 62 12.3.2 Where to get AMD, the automounter daemon. . . . . . . . . . . . . . . . . . . . . . . 63 12.3.3 An example AMD configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 12.4 Using Linux as a router . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 12.5 NIS - Sun Network Information System. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65

13 Experimental and Developmental modules. 65

13.1 AX.25 - A protocol used by Amateur Radio Operators. . . . . . . . . . . . . . . . . . . . . . 65 13.1.1 Where to obtain the AX.25 software. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 13.1.2 Installing the AX.25 software. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 13.1.3 Configuring and using the AX.25 software. . . . . . . . . . . . . . . . . . . . . . . . . 67

13.2 Z8530 SCC driver. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 13.3 Ottawa PI/PI2 card driver. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 13.4 snmp agent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 13.5 Experimental ARCNet driver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 13.6 Experimental Token Ring driver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 13.7 V.35 interface board . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 13.8 IPX bridge program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 13.9 IPX RIP and SAP support. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 13.10Demand Dial SLIP/PPP package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

14 Diagnostic tools - How do I find out what is wrong? 72 14.1 ping - are you there? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 14.2 traceroute - How do I get there? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 14.3 tcpdump - capturing and displaying network activity. . . . . . . . . . . . . . . . . . . . . . . . 74

15 Some Frequently Asked Questions, with brief Answers. 75 15.1 General questions: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 15.2 Error messages: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 15.3 Routing questions: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 15.4 Using Linux with fileservers/NFS: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 15.5 slip questions: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 15.6 PPP questions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

16 Known Bugs. 78

17 Copyright Message. 79 18 Miscellaneous, and Acknowledgements. 79

1 Introduction.

This is the Linux NET-2-HOWTO. This document is a complete rewrite of the earlier NET-FAQ, and of the subsequent NET-2-HOWTO versions 1.0+, for the new NET-2 and NET-3 tcp/ip networking code for Linux kernels 1.0 and above.

1.1 Changes from the previous release.

Additions:
Added mailing list details for ARCNet support.
Added IPXBridge software.
Added Dynamic Address allocation for sliplogin slip server.
I've based all dip information on dip337j-uri since dip has been
   dropped from the net-tools releases, let me know if I've forgotten
   something.
diald - demand dialling has been added to experimental code.
dslip - Matt Dillons slip package.
Added brief descriptions of the kernel Makefile networking options.
Added network diagnostic tools section.
Added sample gated config file.
Added reference to Stevens "Unix Network Programming" book.
Added samples of other config files for completeness.
Added the alpha version of IPXRIPD
Added reference to the PPP-HOWTO.
Added other bits and pieces.
Corrections:
Removed the NIS section and added reference to NIS-HOWTO.
Updated kernel and libc versions.
Updated Token Ring driver information.
Updated ARCNet driver information.
Updated dip ownership information - thanks Paul Lucassen
Ammended location details of NetKit package
Ammended version of sliplogin package.
Updated PPP information to 2.1.2b

1.2 A brief development history of Linux Networking.

Ross Biro <biro@yggdrasil.com> wrote the original kernel based networking code for Linux. He

used ethernet drivers written by Donald Becker <becker@cesdis1.gsfc.nasa.gov>, a slip driver

written by Laurence Culhane <loz@holmes.demon.co.uk>, and a D-Link driver by Bj0rn Ekwall
<bj0rn@blox.se>.

The further development of the Linux networking code was later taken up by Fred van Kempen <waltje@hacktic.nl>, who took Ross’s code and produced the NET-2 release of network code. NET-2 went through a number of revisions until release NET-2d, when Alan Cox <iialan@iifeak.swan.ac.uk> took Fred’s NET-2d code and set about debugging the code with the aim of producing a stable and working release of code for incorporation into the standard kernel releases. This code was called NET-2D(ebugged), and has been incorporated into the standard kernel releases since some time before Linux vers 1.0 was released.

PPP support was added by Michael Callahan, <callahan@maths.ox.ac.uk> and Al Longyear,

<longyear@netcom.com>, originally as patches to the kernel, and in later releases as part of the standard kernel distribution.

With the release of Linux vers 1.0, Linus made a decision to continue supporting Alan’s code as the ‘standard’ network kernel code.

The latest revision of the code, NET-3, appears in kernel releases 1.1.5 and later, and is essentially the same code, but with many fixes, corrections and enhancements.

Alan has added such features as IPX and AX.25 modules. Florian La Roche, <flla@stud.uni-sb.de> has produced an updated distribution of network applications.

Many other people have made contributions by way of bug fixes, ports of applications and by writing device drivers.

2 Disclaimer.

The Linux networking code is a brand new implementation of kernel based tcp/ip networking. It has been developed from scratch, and is not a port of any existing kernel networking code.

Because it is a fresh implementation it may still have a number of bugs or problems with it, and there may be a number of fixes and patches released. If you are worried about problems then just stick to the version of network code released with the standard kernel releases and utility sets. The networking code has a small team of dedicated people working on it, with a cast of thousands testing the code, and collecting and reporting bugs and problems. Any problem you experience is likely to have already been reported, and be being worked on, and will possible be corrected soon, so be patient, or if you can help, offer your assistance.

We do not, and cannot, know everything there is to know about the Linux network software. Please accept and be warned that this document probably does contain errors. Please read any README files that are included with any of the various pieces of software described in this document for more detailed and accurate information. We will attempt to keep this document as error-free and up-to-date as possible. Versions of software are current as at time of writing.

NOTE: While its name may appear similar to the Berkeley Software Distribution NET-2 release, the Linux network code actually has nothing at all to do with it. Please don’t confuse them.

3 Questions already ?

‘The only stupid question is the unasked one.’

If you have general configuration questions, and you have been unable to find the answers after reading the other various HOWTO and FAQ files, then you would be best served to post them to comp.os.linux.help, or, if you believe your question to be specifically related to the Linux Network code, then you could post it to the NET mailing list. Please include as much relevant information as possible, there is nothing more annoying than to have a bug or problem reported without sufficient information to even begin searching for it.

Version numbers and revisions of code, a detailed account of the problem, and the circumstances that caused it to occur, are essential. Trace and debug messages where available should also be considered mandatory.

If you have a question relating to the configuration of, or problems experienced with, any linux distribution, regardless of who has provided it, please contact the prople who created the distribution first, before attempting to report the problem to the network code developers. The reason for this is that some of the distributions use non-standard directory structures, and supply test/non-standard versions of code and utilities. The developers of the NET-2 code cannot be expected to offer support for the network code as distributed in any form, other than as described in this document, or as per distributed Alpha/Beta test instructions.

To join the Linux NET channel on the mail list server, send mail to:

     linux-activists@niksula.hut.fi
     with the line:
     X-Mn-Admin: join NET
     at the top of the message body (not the subject line).

Remember, keep in mind that the NET channel is for development discussions only.

Note also that a PPP list has been established. To join it, use the same procedure as for joining the NET channel, except specify PPP in place of NET in the X-Mn-Admin: field.

Note also that a HAMS list has been established. This list has been established for the discussion of programs related to Amateur Radio. To join it, follow the same procedure as for joining the NET or PPP channels, except specify HAMS in place of NET in the X-Mn-Admin: field.

4 Related Documentation. (Where to learn about tcp/ip)

If you are looking for information about tcp/ip networking that this HOWTO does not cover, then you might try the following sources, as they provide some very useful information.

Olaf Kirch has written a substantial document as part of the Linux Documentation Project entitled the Linux Network Administration Guide. This is an excellent document. It covers all aspects of setting up and using the tcp/ip networking under Linux, including NFS, UUCP, mail, News, nameserver etc.

Olaf’s book supplements this HOWTO, taking up where this document leaves off. This document covers the installation and configuration of the NET code, i.e. ‘How to put your machine on the net’. If you are new to unix networking, then I strongly urge you to obtain a copy and read it first. It will answer a lot of questions for you that are not within the scope of this document.

The current release version is available in:

sunsite.unc.edu

     /pub/Linux/docs/linux-doc-project/network-guide/*

There are various versions of the document in this directory. The most common formats are supported, being plain ascii, Postscript, DVI, Latex and groff.

The Linux Network Administrators Guide is Copyright (c) by Olaf Kirch.

You should also read the other HOWTO documents relevant to networking with Linux.

They are:

The Ethernet-HOWTO (ftp://sunsite.unc.edu/pub/Linux/docs/HOWTO/Ethernet-HOWTO) which you

should read if you intend using an ethernet card with Linux. It includes much more detail on how to select, install and configure an ethernet card for Linux.

The PPP-HOWTO (http://sunsite.unc.edu/mdw/HOWTO/PPP-HOWTO.html) if you intend using ppp.
The Serial-HOWTO (http://sunsite.unc.edu/mdw/HOWTO/Serial-HOWTO.html) if you intend using slip

or ppp in server mode.

The Mail-HOWTO (http://sunsite.unc.edu/mdw/HOWTO/Mail-HOWTO.html) and the News-HOWTO
(http://sunsite.unc.edu/mdw/HOWTO/News-HOWTO.html) for some specific information on setting up Mail

and News on your system.

The UUCP-HOWTO (http://sunsite.unc.edu/mdw/HOWTO/UUCP-HOWTO.html) if you will be connecting

to the net via UUCP.

The NIS-HOWTO (http://sunsite.unc.edu/mdw/HOWTO/NIS-HOWTO.html) if you are interested in run-

ning a version of Sun’s Network Information Service.

For more general information on Unix network configuration another good place to look for help on setting up your network is the O’Reilly and Associates book TCP/IP Network Administration, (the one with the Crab on the cover). Keep in mind that the Linux Network code is now a fairly standard implementation of tcp/ip networking, this means that the commands to configure and use it will work in much the same way as for those for other unix operating systems. Keep in mind though that some of the arguments and options might differ slightly from those in the book.

If you are after some basic tutorial information on tcp/ip networking generally, then I recommend you take a look at the following documents:

tcp/ip introduction

     text version (ftp://athos.rutgers.edu/runet/tcp-ip-intro.doc),  postscript version
     (ftp://athos.rutgers.edu/runet/tcp-ip-intro.ps).

tcp/ip administration

     text version (ftp://athos.rutgers.edu/runet/tcp-ip-admin.doc),  postscript version
     (ftp://athos.rutgers.edu/runet/tcp-ip-admin.ps).

If you are after some more detailed information on tcp/ip networking then I highly recommend:

     "Internetworking with TCP/IP"
     by Douglas E. Comer
     ISBN 0-13-474321-0
     Prentice Hall publications.

If you are wanting to learn about how to write network applications in a Unix compatible environment then I also highly recommend:

     "Unix Network Programming"
     by W. Richard Stevens
     ISBN 0-13-949876-1
     Prentice Hall publications.

4.1 New versions of this document.

If your copy of this document is more than a two months old then I strongly recommend you obtain a newer version. Because the networking support for Linux is changing so rapidly this document also changes fairly frequently. The latest released version of this document can always be retrieved by anonymous ftp from:

sunsite.unc.edu

     /pub/Linux/docs/HOWTO/NET-2-HOWTO
     or:
     /pub/Linux/docs/HOWTO/other-formats/NET-2-HOWTO{-html.tar,ps,dvi}.gz

or  via  the  World  Wide

Web from the Linux Documentation Project Web Server (http://sunsite.unc.edu/mdw/linux.html), at

page: NET-2-HOWTO (http://susnite.unc.edu/mdw/HOWTO/NET-2-HOWTO.html) or directly from me,
<terryd@extro.ucc.su.oz.au>. It will also be posted to the newsgroups: comp.os.linux.announce,
comp.os.linux.help, and news.answers periodically.

You can find news.answers FAQ postings, including this one, archived on rtfm.mit.edu:/pub/usenet.

4.2 Feedback.

Please send any comments, updates, or suggestions to me, <terryd@extro.ucc.su.oz.au>. The sooner I get feedback, the sooner I can update and correct this document. If you find any problems with it, please mail me instead of posting to one of the newsgroups, as I may miss it.

5 Some terms used in this document.

You will often see the terms client and server used in this document. They are normally fairly specific terms but in this document I have generalised their definitions a little so that they mean the following:

client The machine or program that initiates an action or a connection for the purpose of gaining use of some service or data.

server The machine or program that accepts incoming connections from multiple remote machines and provides a service or data to those.

These definitions are not very reliable either, but they provide a means of distinguishing the ends of peer to peer systems such as slip or ppp which truly do not actually have clients and servers.

Other terms you will see are:

datagram A datagram is a discrete package of data and headers which contain addresses, which is the basic unit of transmission across an IP network. You might also hear this called a ‘packet’.

MTU The Maximum Transmission Unit (MTU ) is a parameter that determines the largest datagram than can be transmitted by an IP interface without it needing to be broken down into smaller units. The MTU should be larger than the largest datagram you wish to transmit unfragmented. Note, this only prevents fragmentation locally, some other link in the path may have a smaller MTU and the datagram will be fragmented there. Typical values are 1500 bytes for an ethernet interface, or 576 bytes for a SLIP interface.

MSS The Maximum Segment Size (MSS) is the largest quantity of data that can be transmitted at one time. If you want to prevent local fragmentation MSS would equal MTU-IP header.

window The window is the largest amount of data that the receiving end can accept at a given point in time.

route The route is the path that your datagrams take through the network to reach their destination.

6 NET-2/NET-3 Supported functionality.

The NET code is a complete kernel based implementation of tcp/ip for Linux. The recent NET-2 and NET-3 versions of code support:

Ethernet Cards

Most popular ethernet cards are supported. SLIP (Serial Line IP) and PPP

for tcp/ip networking over serial lines such as the telephone via modem, or a local cable between two machines.

Van Jacobsen Header Compression

for compressing the tcp/ip headers to improve slip/ppp performance over low speed lines. PLIP (Parallel Lines IP)

to allow local connections between two machines using your printer ports. NFS (Networked File System)

to allow you to remotely mount another machines filesystems. AX.25 (A protocol used by Amateur Radio Operators)

Alan Cox has some experimental code working. PI Card (An 8530 SCC based card used by Amateur Radio Operators)

An experimental PI Card driver is available. IPX/SPX (Novell)

to allow you to write custom SPX/IPX applications, or to use Linux as an IPX router. Sun’s Network Information System - NIS

An NIS implementation has been ported to Linux should you wish to use it.

The NET-2 and NET-3 network code does not yet currently support:

NCP (Novell Netware) support to allow Linux to serve and mount Novell network devices. This is being worked on but due to the proprietry nature of the product it may take some time.

Lan types other than ethernet This means token ring, arcnet, FDDI, etc. An experimental Token Ring driver is being developed. An experimental ARCNet driver has also been developed. Details of both are included later.

ISDN Support this is being developed. I would appreciate any more up to date information.

6.1 Supported Ethernet cards.

The standard linux kernel release supports the following type of Ethernet cards:

• NE2000/NE1000 and close compatibles.

• WD80*3 and close compatibles.

• SMC Ultra

• 3c501 (obselete and very slow)

• 3c503 and close compatibles.

• 3c509/3c579

• HP PCLAN (however a newer kernel is required for the HP PCLAN+)

• AT1500 and NE2100 (LANCE and PCnet-ISA) and close compatibles.

• AT1700 (not clones)

• DEPCA and close compatibles.

• D-Link DE600 pocket adaptor and close compatibles.

• AT-LAN-TEC/RealTek pocket adaptor and close compatibles.

Additional drivers are available in the 1.1.* and later development kernels. The Ethernet-HOWTO contains a lot of very useful information on the supported ethernet cards, including information on how to choose an ethernet card if you are intending to puchase some specifically for Linux. As mentioned above, Linux supports other means of network connection if you don’t have access to an ethernet card or connection. Many universities and businesses worldwide offer some form of dial-up network access. Generally these forms of access will offer an option of either SLIP or PPP access, so you will be well catered for. All you will need is a telephone modem, the one you already have may well be good enough, and to configure your Linux system appropriately. There are sections below that describe exactly what you need.

7 Getting the NET-2/NET-3 software.

Before you can configure the networking software you must obtain all of the bits and pieces that make it up. These include the current version of the kernel code (version 1.0 or later), the correct system libraries, the tcp/ip configuration programs and files (e.g. /sbin/ifconfig, /etc/hosts etc.), and finally a set of network application programs (such as telnet, ftp, rlogin etc.). If you obtained Linux from a distribution you may already have all that you need. Check and make sure that you do. For example, some Linux distributions come with all of the network configuration files, binaries, libraries, and kernel installed, so there’s no reason to get the following files.

NOTE: they may be in directories and files different to those specified in this HOWTO document

If you DO have the network software, skip to the ‘Configuring the kernel’ section. If you DO NOT have the network software follow the following directions.

7.1 The kernel source.

Version 1.0 of the Linux kernel is the production version. Any of the Linux kernels after that release are enhancements or bug fixes. If you feel at all concerned about the possibility of having to patch and modify the kernel source, then you should stick to this release, as it will do most of what you want it to. In the case of the networking code though, I strongly suggest you just take a deep breath and follow the newer releases of code, as there have been many changes in the newer version kernels that affect networking. I know you hear it from everyone and everywhere, but when trying out any new version of kernel software you should always ensure that you have sufficient backups of your system just in case something goes seriously wrong while you are testing.

The current kernel version is found in:

ftp.funet.fi

     /pub/OS/Linux/PEOPLE/Linus/v1.1/v1.1.52.tar.gz

This is a gzipped file, so you will need gzip to uncompress it.

To install it, try:

     # cd /usr/src
     # mv linux linux.old
     # gzip -dc v.1.1.52.tar.gz | tar xvf -

You may also find some files called patch46.gz ... in the same directory. These are patch files. If you have a linux kernel that is version 1.1.52 then what this means is that you have linux kernel version 1.1.0 with patches 1 to 52 applied, so you won’t need to apply any of these. If there are any patch files that are greater than the version of kernel you have, you should obtain all of those above, and apply them, in sequence, with something like the following commands:

     # cd /usr/src
     # gzip -dc .../patch1.gz | patch -p0
     # gzip -dc .../patch2.gz | patch -p0
     # gzip -dc .../patch3.gz | patch -p0
      ...

7.2 The libraries.

You’ll want at least version 4.4.2 of libc, as there were problems with earlier version that affected subnet masks.

The current libraries (libc-4.6.20) can be found in: sunsite.unc.edu

     /pub/Linux/GCC/

You will need at least the following files:

• image-4.6.20.tar.gz

• inc-4.6.20.tar.gz

• extra-4.6.20.tar.gz

• release.libc-4.6.20

You MUST read release.libc-4.6.20 before you install the libraries. Please note that to use release 4.5.26 you will also need at least GCC version 2.6.2, and Linux kernel 1.1.52 or later.

7.3 The network configuration tool suite.

You will need the utility suite that provides tools to configure your network support. The current NET-2 utility suite is available from: sunacm.swan.ac.uk

     /pub/misc/Linux/Networking/PROGRAMS/NetTools/net-tools-1.1.56.tar.gz

Because the kernel networking code is still changing some changes to the network tools have been necessary as new kernels are released, so you will need to choose the version that is appropiate for the kernel version you intend to use. The filenames reflect the earliest version of kernel that the tools will work with. Please choose the filename whose version equals, or is less than the version of kernel source you intend to use. To build and install the tools, you should try:

     # cd /usr/src
     # mkdir net-tools
     # cd net-tools
     # gzip -dc net-tools-1.1.56.tar.gz | tar xvf -
     # make config
     # make
     If everything makes ok, then:
     # make install

If you use a kernel version 1.1.26 or earlier you should look in:

sunacm.swan.ac.uk

     /pub/misc/Linux/Networking/PROGRAMS/Other/net032/

In this directory you will find three versions of the network tools. The following table lists net-032 package name with the relevant kernel versions:

     net-0.32d-net3.tar.gz     1.1.12+
     net-0.32b.tar.gz          1.1.4+
     net-0.32.old.tar.gz       pre 1.1.4 kernels

These packages include the essential network configuration programs such as ifconfig, route, netstat etc. These will be discussed later.

7.4 The network applications.

You will want a number of network application programs. These are programs like telnet, ftp, finger and their daemons at least. Florian La Roche, <flla@stud.uni-sb.de> has put together a fairly complete distribution of network applications in both binary and source form. The tcp/ip application binaries and some sample config files are found in:

ftp.funet.fi

     /pub/OS/Linux/PEOPLE/Linus/net-source/base/NetKit-A-0.07.bin.tar.gz
     /pub/OS/Linux/PEOPLE/Linus/net-source/base/NetKit-B-0.06.bin.tar.gz

If there are newer versions then use the newer versions. Please read the README file first just to make sure that you have the necessary prerequisites.

Be sure to make backups of any config files important to you. If this is a new installation you probably don’t have any. You can unpack each of the packages above with the command:

     # cd /
     # gzip -dc filename.tar.z | tar xpvlf -

IMPORTANT NOTE: Florian has built and prepackaged these tar files for your convenience. Florian has attempted to make them as complete as possible and has included a distribution of the binaries found in the net-tools-n.n.nn releases. Unfortunately Florian has chosen not to use the same directory structure as Alan did when he prepared the installation script for the net-tools. This will mean that you should be very careful when installing them. Florian will change this later so that this difference is not a problem, but until then, I suggest you do the following instead of the above:

     -  Unpack the binaries somewhere safe:
     # cd /usr/src
     # mkdir NetKit

     # cd NetKit
     # gzip -dc NetKit-A-0.07.bin.tar.gz | tar xpvlf -
     # gzip -dc NetKit-B-0.06.bin.tar.gz | tar xpvlf -
     -  Remove Florians copies of the network tools previously described:
     # rm ./bin/hostname ./sbin/route ./sbin/ifconfig ./sbin/netstat
     # rm ./usr/sbin/arp ./usr/sbin/rarp ./usr/sbin/slattach
     -  Copy Florian's files into their new home:
     # cp -vrpd . /

7.5 Additional drivers or packages.

If you want to add some developmental, or Alpha/Beta test code, such as AX.25 support, you will need to obtain the appropriate support software for those packages. Please check the relevant sections for those packages in this document for more detail.

8 Configuring the kernel.

Before you can use any of the network tools, or configure any network devices, you must ensure that your kernel has the necessary network support built into it. The best way of doing this is to compile your own, selecting which options you want and which you don’t.

Assuming you have obtained and untarred the kernel source already, and applied any patches that you might need to have applied to get any nonstandard or developmental software installed, all you have to do is edit /usr/src/linux/drivers/net/CONFIG. This file has many comments to guide you in editing it, and in general you will need to edit very little, as it has sensible defaults. In my case I don’t need to edit it at all. This file is really necesary if your ethernet card is an unusual one, or is one that isn’t automatically detected by the ethernet driver. It allows you to hard code some of the elements of your ethernet hardware. For example, if your ethernet card is a close, but not exact clone of a WD-8013, then you might have to configure the shared memory address to ensure the driver detects and drives the card properly. Please check the Ethernet-HOWTO for more definitive information on this file and its effect on ethernet cards. This file also contains configurable parameters for PLIP, though the defaults should again be ok unless you have a particularly slow machine.

When you are happy that the CONFIG file is suitable for your purposes, then you can proceed to build the kernel. Your first step will be to edit the top level Makefile to ensure the kernel will be built with the appropriate VGA settings, and then you must run the kernel configuration program:

     # cd /usr/src/linux
     # make config

You will be asked a series of questions. There are four sections relevant to the networking code. They are

the General setup, Networking options, Network device support, and the Filesystems sections. The

most difficult to configure is the Network device support section, as it is where you select what types of 8. Configuring the kernel. 17

physical devices you want configured. On the whole you can just use the default values for the other sections fairly safely. The following will give you an idea of how to proceed:

*
* General setup
*
 ...
 ...
Networking support (CONFIG_NET) [y] y
 ...
 ...

In the General setup section you simply select whether you want network support or not. Naturally you must answer yes.

*
* Networking options
*
TCP/IP networking (CONFIG_INET) [y] y
IP forwarding/gatewaying (CONFIG_IP_FORWARD) [y] y
*
* (it is safe to leave these untouched)
*
PC/TCP compatibility mode (CONFIG_INET_PCTCP) [n] n
Reverse ARP (CONFIG_INET_RARP) [n] n
Assume subnets are local (CONFIG_INET_SNARL) [y] y
Disable NAGLE algorithm (normally enabled) (CONFIG_TCP_NAGLE_OFF) [n] n
The IPX protocol (CONFIG_IPX) [n] n

The second half of the Networking options section allows you to enable or disable some funky features that you can safely accept the defaults on until you have some idea why you want to change them. They are described briefly later if you are interested.

*
* Network device support
*
Network device support? (CONFIG_NETDEVICES) [y] y
Dummy net driver support (CONFIG_DUMMY) [n] n
SLIP (serial line) support (CONFIG_SLIP) [y] y
 CSLIP compressed headers (SL_COMPRESSED) [y] y
PPP (point-to-point) support (CONFIG_PPP) [y] y
Load balancing support (experimental) (CONFIG_SLAVE_BALANCING) [n] n
Do you want to be offered ALPHA test drivers (CONFIG_NET_ALPHA) [n] n
Western Digital/SMC cards (CONFIG_NET_VENDOR_SMC) [y] y
WD80*3 support (CONFIG_WD80x3) [y] y
SMC Ultra support (CONFIG_ULTRA) [n] n
3COM cards (CONFIG_NET_VENDOR_3COM) [n] n

     Other ISA cards (CONFIG_NET_ISA) [n] n
     PLIP (parallel port) support (CONFIG_PLIP) [n] n
     EISA and on board controllers (CONFIG_NET_EISA) [n] n
     Apricot Xen-II on board ethernet (CONFIG_APRICOT) [n] n
     Pocket and portable adaptors (CONFIG_NET_POCKET) [n] n
     *

This section is the most important, and the most involved. It is where you select what hardware devices you want to support. You can see that I have selected SLIP support with header compression, PPP, the WD80*3 driver, and nothing else. Simply answer `y' to whatever you want to play with, and `n' to that you don’t.

     *
     * Filesystems
     *
      ...
      ...
     /proc filesystem support (CONFIG_PROC_FS) [y]
     NFS filesystem support (CONFIG_NFS_FS) [y]
      ...
      ...

If you wish to run an NFS client then you will want to include the NFS filesystem type. You will need to include the /proc filesystem because a number of the network utilities use it.

After you have completed the configuration, all that remains is to actually compile the kernel:

     # make dep
     # make

Don’t forget to make zlilo if the new kernel compiles and tests ok.

8.1 What do all those funky Networking options actually do?

Newer kernels have a number of options that you are asked about when you do a make config. Generally you will not need to change these, but some of the options might be useful to you in certain circumstances.

TCP/IP networking

This one is obvious, it selects whether you configure the tcp/ip suite of protocols into your kernel. Chances are if you are reading this then you will want to answer ‘y’ to this one.

Dummy networking device

This was added to allow slip and PPP users to configure an address on their linux machine that would not be dependent on their serial link being established. It is an easy way to give your linux machine two addresses.

IP forwarding/gatewaying This determines what your kernel will do when it receives a datagram that has a destination address that is not one of its own devices. You must have this option selected if you want your kernel to act as an IP router. Most SLIP and PPP servers will want this option selected.

IP multicasting This is alpha test code support for IP multicasting, examples of which include services such as ‘Internet Talk Radio’ and live video. You will need additional programs to make use of this facility, this is just the kernel support.

IP firewalling This option allows you to provide flexible security options for your linux machine. You can selectively enable/disable access to tcp/ip ports from any address ranges you choose. This also needs additional programs to support it.

IP accounting This option is for those people that want to use their Linux machine to provide internet connectivity to others on a commercial basis. It allows you to count and record incoming and outgoing bytes on a per port and address basis. With the addition of suitable software this would allow you to produce seperate usage charges for each person using your systems networking capabilities.

PC/TCP compatibility mode This option provides a work-around for a bug that causes problems when using the PC/TCP networking programs to talk to your linux machine. There is a PC/TCP bug which provokes a difficult to remedy Linux bug, and this option prevents the two clashing. Normally you would leave this disabled, but if you have users on your network who use PC/TCP then you may have to enable this option to prevent problems.

Reverse ARP This option allows you to configure the RARP protocol into your kernel. This option was added to allow the booting of Sun 3 systems. This is not generally very useful otherwise.

Assume subnets are local This option selects whether you assume that your whole subnet is directly connected to your linux machine, or whether it might be bridged or otherwise subdivided at a lower layer. In practise it will make little difference if you leave it set at the default.

Disable NAGLE algorithm This is a timing option that determines when a datagram should be transmitted. The default setting provides for the best throughput in most situations and you should leave this set as it is, as disabling it will degrade your throughput. This option can be selectively changed from within a program with a socket option, and you would normally be much better off leaving it set at the default and specifically writing your programs to disable the NAGLE algorithm if they require extremely fast interactivity.

The IPX protocol This option selects whether you compile the IPX protocol support into your kernel. The IPX protocol is an internetworking protocol similar in function to the IP protocol. This protocol is one of those used by the Novell suite.

Amateur Radio AX.25 Level 2 This option selects whether you compile in the Amateur Radio AX.25 protocol suite. If you select this option then a new class of network sockets are available for programming. The AX.25 protocol is used primarily by Amateur Radio Operators for packet radio use.

9 Configuring the Network Devices.

If everything has gone ok so far, then you will have a Linux kernel which supports the network devices you intend to use, and you also have the network tools with which to configure them. Now comes the fun part! You’ll need to configure each of the devices you intend to use. This configuration generally amounts to telling each device things like what its IP address will be, and what network it is connected to.

In past versions of this document I have presented near complete versions of the various configuration files and included comments to modify or delete lines from them as appropriate. From this version onwards I will take a slightly different approach which I hope will result in you having a complete set of uncluttered configuration files that you have built from scratch so you know exactly what is in them, and why. I’ll describe each of these files, and their function, as we come to them.

9.1 Configuring the special device files in /dev

You do not need to configure any special device files in the /dev directory for Linux Networking. Linux does not need or use them as other operating systems might. The devices are built dynamically in memory by the kernel, and since they are only names there is no need for them to have an appearance directly to you. The kernel provides all of the programming hooks and interfaces that you need to utilise them effectively.

9.2 What information do I need before I begin ?

Before you can configure the networking software, you will need to know a number of pieces of information about your network connection. Your network provider or administrator will be able to provide you with most of them.

9.2.1 IP Address.

This is the unique machine address, in dotted decimal notation, that your machine will use. An example is 128.253.153.54. Your network administrator will provide you with this information.

If you will be using a slip or plip connection you may not need this information, so skip it until we get to the slip device.

If you’re using the loopback device only, ie no ethernet, slip or plip support, then you won’t need an ip address as the loopback port always uses the address 127.0.0.1.

9.2.2 Network Mask (‘netmask’).

For performance reasons it is desirable to limit the number of hosts on any particular segment of a network. For this reason it is common for network administrators to divide their network into a number of smaller networks, known as subnets, which each have a portion of the network addresses assigned to them. The network mask is a pattern of bits, which when overlayed onto an address on your network, will tell you which subnetwork it belongs to. This is very important for routing, and if you find for example, that you can happily talk to people outside your network, but not to some people on your own network, then it is quite likely that you have specified an incorrect subnet mask.

Your network adminstrators will have chosen the netmask when the network was designed, and therefore they should be able to supply you with the correct mask to use. Most networks are class-C subnetworks which use 255.255.255.0 as their netmask. Other larger networks use class-B netmasks (255.255.0.0). The NET-2/NET-3 code will automatically select a default mask when you assign an address to a device. The default assumes that your network has not been subnetted.

The NET-2/NET-3 code will choose the following masks by default:

     For addresses with the first byte:
     1-127         255.0.0.0         (Class A)
     128-191       255.255.0.0       (Class B)
     192+          255.255.255.0     (Class C)

if one of these doesn’t work for you, try another. If this doesn’t work ask your network administrator or local network guru (dime a dozen) for help.

You don’t need to worry about a netmask for the loopback port, or if you are running slip/plip.

9.2.3 Network Address.

This is your IP address masked (bitwise AND) with your netmask. For example:

     If your netmask is:           255.255.255.0
     and your IP address is:       128.253.154.32    &&
                                   ---------------
     your Network address is:      128.253.154.0     =

9.2.4 Broadcast Address.

‘A shout is a whisper that everyone hears whether they need to or not’

This is normally your network address logically ORed with your netmask inverted. This is simpler than it sounds. For a Class-C network, with network mask 255.255.255.0, your Broadcast Address will be your network address (calculated above), logically ORed with 0.0.0.255, the network mask inverted.

A worked example might look like:

     If your netmask is:          255.255.255.0      !
     the netmask inverted is:       0.  0.  0.255    =
     If your Network address is:  128.253.154.0      ||
                                  ----------------
     Your broadcast address is:   128.253.154.255    =

Note that for historical reasons some networks use the network address as the broadcast address. If you have any doubts contact your network administrator.

If you have access to a sniffer, or some other device capable of providing you with a trace of your network traffic, then you might be able to determine both the network and broadcast addresses by watching other traffic on the lan. Keep an eye open for, (or filter everything except), ethernet frames destined for the ethernet broadcast address: ff:ff:ff:ff:ff:ff. If any of them has an IP source address of your local router, and the protocol ID is not ARP, then check the destination IP address, because this datagram may well be a RIP routing broadcast from your router, in which case the destination IP address will be your broadcast address.

Once again, if you’re not sure, check with your network administrator, they’d rather help you, than have you connect your machine misconfigured.

9.2.5 Router (‘Gateway’) Address.

‘There must be some way out of here.’

This is the address of the machine that connects your network to the rest of the Internet. It is your ‘gateway’ to the outside world. A couple of conventions exist for allocating addresses to routers which your network might follow, they are: The router is the lowest numbered address on the network, the router is the highest numbered host on the network. Probably the most common is the first, where the router will have an address that is mostly the same as your own, except with a .1 as the last byte. eg. if your address is 128.253.154.32, then your router might be 128.253.154.1. The router can in fact have any address valid on your network and function properly, the address doesn’t matter at all. There may in fact even be more than one router on your network. You will probably need to talk to your network adminstrator to properly identify your router address.

If you’re using only loopback then you don’t need a router address. If you’re using PPP then you also don’t need your router address, because PPP will automatically determine the correct address for you. If you’re using SLIP, then your router address will be your SLIP server address.

9.2.6 Nameserver Address.

Most machines on the net have access to a name server which translates human tolerable hostnames into machine tolerable addresses, and vice versa. Your network administrators will again tell you the address of your nearest nameserver. You can in fact run a nameserver on your own machine by running named, in which case your nameserver address will be 127.0.0.1, the loopback port address. However it is not required that you run named at all; see section ‘named’ for more information.

If you’re only using loopback then you don’t need to know the nameserver address since you’re only going to be talking to your own machine.

9.2.7 NOTE for SLIP/PLIP/PPP users.

You may or may not in fact need to know any of the above information. Whether you do or not will depend on exactly how your network connection is achieved, and the capabilities of the machine at the other end of the link. You’ll find more detail in the section relevant to configuration of the SLIP/PLIP and PPP devices.

9.3 /etc/rc.d/rc.inet1,2 or /etc/rc.net

While the commands to configure your network devices can be typed manually each time, you will probably want to record them somewhere so that your network is configured automatically when you boot your machine.

The `rc' files are specifically designed for this purpose. For the non-unix-wizard: `rc' file are run at bootup time by the init program and start up all of the basic system programs such as syslog, update, and cron. They are analagous to the MS-DOS autoexec.bat file, and rc might stand for ‘runtime commands’. By convention these files are kept under the /etc directory. The Linux Filesystem Standard doesn’t go so far as to describe exactly where your rc files should go, stating that it is ok for them to follow either the BSD (/etc/rc.*) or System-V (/etc/rc.d/rc*) conventions. Alan, Fred and I all use the System-V convention, so that is what you will see described here. This means that these files are found in /etc/rc.d and are called rc.inet1 and rc.inet2 . The first rc file that gets called at bootup time is /etc/rc, and it in turn calls others, such as rc.inet1, which in turn might called rc.inet2. It doesn’t really matter where they are kept, or what they are called, so long as init can find them.

In some distributions the rc file for the network is called rc.net and is in the /etc subdirectory. The rc.net file on these systems is simply the rc.inet1 and the rc.inet2 files combined into one file that gets executed. It doesn’t matter where the commands appear, so long as you configure the interfaces before starting the network daemons and applications.

I will refer to these files as rc.inet1 and rc.inet2 , and I keep them in the /etc/rc.d, so if you are using one of the distributions that uses rc.net, or you want to keep the files somewhere else, then you will have to make appropriate adjustments as you go.

We will be building these files from scratch as we go.

9.3.1 rc.inet1

The rc.inet1 file configures the basic tcp/ip interaces for your machine using two programs: /sbin/ifconfig, and /sbin/route.

ifconfig /sbin/ifconfig is used for configuring your interfaces with the parameters that they require to function, such as their IP address, network mask, broadcast addresses and similar. You can use the ifconfig command with no parameters to display the configuration of all network devices. Please check the ifconfig man page for more detail on its use.

route /sbin/route is used to create, modify, and delete entries in a table (the routing table) that the networking code will look at when it has a datagram that it needs to transmit. The routing table lists destination address, and the interface that that address is reachable via. You can use the route command with no parameters to display the contents of the routing table. Please check the route man page for more detail on its use.

9.3.2 rc.inet2

The rc.inet2 file starts any network daemons such as inetd, portmapper and so on. This will be covered in more detail in section ‘rc.inet2’, so for the moment we will concentrate on rc.inet1 . I have mentioned this file here so that if you have some other configuration, such as a single rc.net file you will understand what the second half of it represents. it is important to remember that you must start your network applications and daemons after you have configured your network devices.

9.4 Configuring the Loopback device (mandatory).

The loopback device isn’t really a hardware device. It is a software construct that looks like a physical interface. Its function is to happily allow you to connect to yourself, and to test network software without actually having to be connected to a network of any kind. This is great if you are developing network software and you have a slip connection. You can write and test the code locally, and then when you are ready to test it on a live network, eatablish your slip connection and test it out. You won’t hurt others users if your program misbehaves.

By convention, the loopback device always has an IP address of 127.0.0.1 and so you will use this address when configuring it.

The loopback device for Linux is called ‘lo’. You will now make the first entry into your rc.inet1 file. The following code fragment will work for you:

     #!/bin/sh
     #
     # rc.inet1   --  configures network devices.
     #
     # Attach the loopback device.
     /sbin/ifconfig lo 127.0.0.1
     #
     # Add a route to point to the loopback device.
     /sbin/route add 127.0.0.1
     # End loopback
     #

You have used the ifconfig program to give the loopback interface its IP address, and route program to create an entry in the routing table that will ensure that all datagrams destined for 127.0.0.1 will be sent to the loopback port.

There are two important points to note here.

Firstly, the netmask and broadcast addresses have been allowed to take the default values for the loopback device described earlier in section ‘Network Mask’. To see what they are, try the ifconfig program without any arguments.

     # ifconfig
     lo        Link encap Local Loopback
               inet addr 127.0.0.1  Bcast 127.255.255.255  Mask 255.0.0.0
               UP BROADCAST LOOPBACK RUNNING  MTU 2000  Metric 1
               RX packets 0 errors 0 dropped 0 overrun 0
               TX packets 30 errors 0 dropped 0 overrun 0
     #

Secondly, its not obvious how the route command chose the loopback device as the device for the route to 127.0.0.1. The route program is smart enough to know that 127.0.0.1 belongs to the network supported by the loopback device. It works this out by checking the IP address and the netmask. You can use the route command with no arguments to display the contents of the routing table:

     # route
     Kernel routing table
     Destination     Gateway         Genmask         Flags Metric Ref    Use Iface
     127.0.0.0       *               255.0.0.0       U     0      0       30 lo
     #

Note: You might want to use the -n argument if your name resolver is not yet configured properly. The -n argument tells route to just display the numeric addresses, and to not bother looking up the name.

9.5 Configuring an ethernet device. (optional)

You’ll only be interested in this section if you wish to configure an ethernet card, if not then skip on ahead to the next section.

To configure an ethernet card is only slightly more complicated than configuring the loopback device. This time you should probably specify explicitly the network mask and the broadcast address, unless you are sure that the defaults will work ok, and they probably will.

For this you will need the IP address that you have been assigned, the network mask in use on your network, and the broadcast address in use.

The first ethernet device for a Linux system is called ‘eth0’, the second ‘eth1’ and so forth. You will now add a section to your rc.inet1 file. The following code fragment will work for you if you change the addresses specified for real ones:

     #

     # Attach an ethernet device
     #
     #  configure the IP address, netmask and broadcast address.
     /sbin/ifconfig eth0 IPA.IPA.IPA.IPA
     /sbin/ifconfig eth0 netmask NMK.NMK.NMK.NMK
     /sbin/ifconfig eth0 broadcast BCA.BCA.BCA.BCA
     #
     # add a network route to point to it:
     /sbin/route add -net NWA.NWA.NWA.NWA device eth0
     #
     # End ethernet
     #

Where:

IPA.IPA.IPA.IPA

represents your IP Address.

NMK.NMK.NMK.NMK

represents your netmask.

BCA.BCA.BCA.BCA

represents your Broadcast address.

NWA.NWA.NWA.NWA

represents your Network Address.

Note the use of the -net argument to the route command. This tells route that the route to be added is a route to a network, and not to a host. There is an alternative method of achieving this, you can leave off the -net if you have the network address listed in the /etc/networks file. This is covered later in section ‘/etc/networks’.

9.6 Configuring a SLIP device (optional)

SLIP (Serial Line Internet Protocol) allows you to use tcp/ip over a serial line, be that a phone line with a dialup modem, or a leased line of some sort. Of course to use slip you need access to a slip-server in your area. Many universities and businesses provide slip access all over the world.

Slip uses the serial ports on your machine to carry IP datagrams. To do this it must take control of the serial device. Slip device names are named sl0 , sl1 etc. How do these correspond to your serial devices ? The networking code uses what is called an ioctl (i/o control) call to change the serial devices into slip devices. There are two programs supplied that can do this, they are called dip and slattach 9.6.1 dip

dip (Dialup IP) is a smart program that is able to set the speed of the serial device, command your modem to dial the remote end of the link, automatically log you into the remote server, search for messages sent to you by the server, and extract information for them such as your IP address, and perform the ioctl necessary to switch your serial port into slip mode. dip has a powerful scripting ability, and it is this that you can exploit to automate your logon procedure.

dip used to be supplied with the net-tools, but since development of dip is now seperate, you have to source it seperately. There have been a number of other versions of dip produced which offer a variety of new features. The dip-uri version seems to be the more popular, but I suggest you take a close look at each to determine which offers enhancements that you find useful. Since dip-uri is is so popular, the examples described in this document are based on current versions of it.

You can find it at:

sunsite.unc.edu

     /pub/Linux/system/Network/serial/dip337j-uri.tgz

To install it, try the following:

     #
     # cd /usr/src
     # gzip -dc dip337j-uri.tgz | tar xvf -
     # cd dip.3.3.7j
      <edit Makefile>
     # make install
     #

The Makefile assumes the existence of a group called uucp, but you might like to change this to either dip or slip depending on your configuration.

9.6.2 slattach

slattach as contrasted with dip is a very simple program, that is very easy to use, but does not have the sophistication of dip. It does not have the scripting ability, all it does is configure your serial device as a slip device. It assumes you have all the information you need and the serial line is established before you invoke it. slattach is ideal to use where you have a permanent connection to your server, such as a physical cable, or a leased line.

9.6.3 When do I use which ?

You would use dip when your link to the machine that is your slip server is a dialup modem, or some other termporary link. You would use slattach when you have a leased line, perhaps a cable, between your machine and the server, and there is no special action needed to get the link working. See section ‘Permanent Slip connection’ for more information.

Configuring slip is much like configuring an Ethernet interface (read section ‘Configuring an ethernet device’ above). However there are a few key differences.

First of all, slip links are unlike ethernet networks in that there is only ever two hosts on the network, one at each end of the link. Unlike an ethernet that is available for use as soon are you are cabled, with slip, depending on the type of link you have, you may have to initialise your network connection in some special way.

If you are using dip then this would not normally be done at boot time, but at some time later, when you were ready to use the link. It is possible to automate this procedure. If you are using slattach then you will probably want to add a section to your rc.inet1 file. This will be described soon.

There are two major types of slip servers: Dynamic IP address servers and static IP address servers. Almost every slip server will prompt you to login using a username and password when dialing in. dip can handle logging you in automatically.

9.6.4 Static slip server with a dialup line and DIP.

A static slip server in one in which you have been supplied an IP address that is exclusively yours. Each time you connect to the server, you will configure your slip port with that address. The static slip server will answer your modem call, possibly prompt you for a username and password, and then route any datagrams destined for your address to you via that connection. If you have a static server, then you may want to put entries for your hostname and IP address (since you know what it will be) into your /etc/hosts. You should also configure some other files such as: rc.inet2, host.conf, resolv.conf, /etc/HOSTNAME, and rc.local. Remember that when configuring rc.inet1, you don’t need to add any special commands for your slip connection since it is dip that does all of the hard work for you in configuring your interface. You will need to give dip the appropriate information, and it will configure the interface for you after commanding the modem to establish the call, and logging you into your slip server.

If this is how your slip server works then you can move to section ‘Using Dip’ to learn how to configure dip appropriately.

9.6.5 Dynamic slip server with a dialup line and DIP.

A dynamic slip server is one which allocates you an IP address randomly, from a pool of addresses, each time you logon. This means that there is no guarantee that you will have any particular address each time, and that address may well be used by someone else after you have logged off. The netework administrator who configured the slip server will have assigned a pool of address for the slip server to use, when the server receives a new incoming call, it finds the first unused address, guides the caller through the login process, and then prints a welcome message that contains the IP address it has allocated, and will proceed to use that IP address for the duration of that call.

Configuring for this type of server is similar to configuring for a static server, except that you must add a step where you obtain the IP address that the server has allocated for you and configure your slip device with that.

Again, dip does the hard work, and new versions are smart enough to not only log you in, but to also be able to automatically read the IP address printed in the welcome message, and store it so that you can have it configure your slip device with it.

If this is how your slip server works then you can move to section ‘Using Dip’ to learn how to configure dip appropriately.

9.6.6 Using DIP.

As explained earlier, dip is a powerful program that can simplify and automate the process of dialling into the slip server, logging you in, starting the connection, and configuring your slip devices with the appropriate ifconfig and route commands.

Essentially to use dip you’ll write a ‘dip script’, which is basically a list of commands that dip understands that tell dip how to perform each of the actions you want it to perform. See sample.dip that comes supplied with dip to get an idea of how it works. dip is quite a powerful program, with many options. Instead of going into all of them here you should looks at the man page, README and sample files that will have come with your version of dip.

You may notice that the sample.dip script assumes that you’re using a static slip server, so you know what your IP address is beforehand. For dynamic slip servers, the newer versions of dip include a command you can use to automatically read and configure your slip device with the IP address that the dynamic server allocates for you. The following sample is a modified verson of the sample.dip that came supplied with dip337j-uri.tgz, and is probably a good starting point for you. You might like to save it as /etc/dipscript and edit it to suit your configuration:

     #
     # sample.dip    Dialup IP connection support program.
     #
     #               This file (should show) shows how to use the DIP
     #       This file should work for Annex type dynamic servers, if you
     #       use a static address server then use the sample.dip file that
     #       comes as part of the dip337-uri.tgz package.
     #
     #
     # Version:      @(#)sample.dip  1.40    07/20/93
     #
     # Author:       Fred N. van Kempen, <waltje@uWalt.NL.Mugnet.ORG>
     #
     main:
       # Next, set up the other side's name and address.
       # My dialin machine is called 'xs4all.hacktic.nl' (== 193.78.33.42)
       get $remote xs4all.hacktic.nl
       # Set netmask on sl0 to 255.255.255.0
       netmask 255.255.255.0

9. Configuring the Network Devices. 30

# Set the desired serial port and speed.
port cua02
speed 38400
# Reset the modem and terminal line.
# This seems to cause trouble for some people!
reset
# Note! "Standard" pre-defined "errlevel" values:
#  0 - OK
#  1 - CONNECT
#  2 - ERROR
#
# You can change those grep'ping for "addchat()" in *.c...
# Prepare for dialing.
send ATQ0V1E1X4\r
wait OK 2
if $errlvl != 0 goto modem_trouble
dial 555-1234567
if $errlvl != 1 goto modem_trouble
# We are connected.  Login to the system.
login:
sleep 2
wait ogin: 20
if $errlvl != 0 goto login_error
send MYLOGIN\n
wait ord: 20
if $errlvl != 0 goto password_error
send MYPASSWD\n
loggedin:
# We are now logged in.
wait SOMEPROMPT 30
if $errlvl != 0 goto prompt_error
# Command the server into SLIP mode
send slip\n
wait SLIP 30
if $errlvl != 0 goto prompt_error
# Get and Set your IP address from the server.
#   Here we assume that after commanding the slip server into SLIP
#   mode that it prints your IP address
get $locip remote 30
if $errlvl != 0 goto prompt_error

       # Set up the SLIP operating parameters.
       get $mtu 296
       # Ensure "route add -net default xs4all.hacktic.nl" will be done
       default
       # Say hello and fire up!
     done:
       print CONNECTED $locip ---> $rmtip
       mode CSLIP
       goto exit
     prompt_error:
       print TIME-OUT waiting for SLIPlogin to fire up...
       goto error
     login_trouble:
       print Trouble waiting for the Login: prompt...
       goto error
     password:error:
       print Trouble waiting for the Password: prompt...
       goto error
     modem_trouble:
       print Trouble ocurred with the modem...
     error:
       print CONNECT FAILED to $remote
       quit
     exit:
       exit

The above example assumes you are calling a dynamic slip server, if you are calling a static slip server, then the sample.dip file that comes with dip337j-uri.tgz should work for you.

When dip is given the get $local command it searches the incoming text from the remote end for a string that looks like an IP address, ie strings numbers seperated by ‘.’ characters. This modification was put in place specifically for dynamic slip servers, so that the process of reading the IP address granted by the server could be automated.

The example above will automaticaly create a default route via your slip link, if this is not what you want, you might have an ethernet connection that should be your default route, then remove the default command from the script. After this script has finished running, if you do an ifconfig command, you will see that you have a device sl0 . This is your slip device. Should you need to, you can modify its configuration manually, after the dip command has finished, using the ifconfig and route commands.

Please note that dip allows you to select a number of different protocols to use with the mode command, the most common example is cslip for slip with compression. Please note that both ends of the link must agree, so you should ensure that whatever you select agrees with what your server is set to.

The above example is fairly robust and should cope with most errors. Please refer to the dip man page for more information. Naturally you could, for example, code the script to do such things as redial the server if it doesn’t get a connection within a prescribed period of time, or even try a series of servers if you have access to more than one.

9.6.7 Permament slip connection using a leased line and slattach.

If you have a cable between two machines, or are fortunate enough to have a leased line, or some other permanent serial connection between your machine and another, then you don’t need to go to all the trouble of using dip to set up your serial link. slattach is a very simple to use utility that will allow you just enough functionality to configure your connection.

Since your connection will be a permanent one, you will want to add some commands to your rc.inet1 file. In essence all you need to do for a permament connection is ensure that you configure the serial device to the correct speed and switch the serial device into slip mode. slattach allows you to do this with one command. Add the following to your rc.inet1 file:

     #
     # Attach a leased line static slip connection
     #
     #  configure /dev/cua0 for 19.2kbps and cslip
     /sbin/slattach -p cslip -s 19200 /dev/cua0 &
     /sbin/ifconfig sl0 IPA.IPA.IPA.IPA pointopoint IPR.IPR.IPR.IPR up
     #
     # End static slip.

Where:

IPA.IPA.IPA.IPA represents your IP address.

IPR.IPR.IPR.IPR represents the IP address of the remote end.

slattach allocated the first unallocated slip device to the serial device specified. slattach starts with sl0 . Therefore the first slattach command attaches slip device sl0 to the serial device specified, and sl1 the next time, etc.

slattach allows you to configure a number of different protocols with the -p argument. In your case you will use either slip or cslip depending on whether you want to use compression or not. Note: both ends must agree on whether you want compression or not.

9.7 Configuring a PLIP device. (optional)

plip (Parallel Line IP), is like slip, in that it is used for providing a point to point network connection between two machines, except that it is designed to use the parallel printer ports on your machine instead of the serial ports. Because it is possible to transfer more than one bit at a time with a parallel port, it is possible to attain higher speeds with the plip interface than with a standard serial device. In addition, even the simplest of parallel ports, printer ports, can be used, in lieu of you having to purchase comparitively expensive 16550AFN UART’s for your serial ports.

Please note that some laptops use chipsets that will not work with PLIP because they do not allow some combinations of signals that PLIP relies on, that printers don’t use.

The Linux plip interface is compatible with the Crywyr Packet Driver PLIP, and this will mean that you can connect your Linux machine to a DOS machine running any other sort of tcp/ip software via plip.

When compiling the kernel, there is only one file that might need to be looked at to configure plip. That file is /usr/src/linux/driver/net/CONFIG, and it contains plip timers in mS. The defaults are probably ok in most cases. You will probably need to increase them if you have an especially slow computer, in which case the timers to increase are actually on the other computer.

To configure a plip interface, you will need to add the following lines to your rc.inet1 file:

     #
     # Attach a PLIP interface
     #
     #  configure first parallel port as a plip device
     /sbin/ifconfig plip0 IPA.IPA.IPA.IPA pointopoint IPR.IPR.IPR.IPR up
     #
     # End plip

Where:

IPA.IPA.IPA.IPA represents your IP address.

IPR.IPR.IPR.IPR represents the IP address of the remote machine.

The pointopoint parameter has the same meaning as for slip, in that it specifies the address of the machine at the other end of the link.

In almost all respects you can treat a plip interface as though it were a slip interface, except that neither dip nor slattach need be, nor can be, used.

9.7.1 PLIP cabling diagram.

plip has been designed to use cables with the same pinout as those commonly used by the better known of the MS-DOS based pc-pc file transfer programs.

The pinout diagram (taken from /usr/src/linux/drivers/net/plip.c) looks as follows:

     Pin Name    Connect pin - pin

     ---------   -------------------------------
     GROUND      25 - 25
     D0->ERROR   2 - 15
     ERROR->D0   15 - 2
     D1->SLCT    3 - 13
     SLCT->D1    13 - 3
     D2->PAPOUT  4 - 12
     PAPOUT->D2  12 - 4
     D3->ACK     5 - 10
     ACK->D3     10 - 5
     D4->BUSY    6 - 11
     BUSY->D4    11 - 6
     D5          7*
     D6          8*
     D7          9*
     STROBE      1*
     FEED        14*
     INIT        16*
     SLCTIN      17*

Notes: Do not connect the pins marked with an asterisk ‘*’. Extra grounds are 18,19,20,21,22,23, and 24.

If the cable you are using has a metallic shield, it should be connected to the metallic DB-25 shell at one end only.

Warning: A miswired PLIP cable can destroy your controller card. Be very careful, and double check every connection to ensure you don’t cause yourself any unnecessary work or heartache.

While you may be able to run PLIP cables for long distances, you should avoid it if you can. The specifications for the cable allow for a cable length of about 1 metre or so. Please be very careful when running long plip cables as sources of strong electromagnetic fields such as lightning, power lines, and radio transmitters can interfere with and sometimes even damage your controller. If you really want to connect two of your computers over a large distance you really should be looking at obtaining a pair of thin-net ethernet cards and running some coaxial cable.

10 Routing. (mandatory)

After you have configured all of your network devices you need to think about how your machine is going to route IP datagrams. If you have only one network device configured then your choice is easy, as all datagrams for any machine other than yours must go via that interface. If you have more than one network interface then your choice is a little more complicated. You might have both an ethernet device and slip connection to your machine at home. In this situation you must direct all datagrams for your machine at home via your slip interface, and all else via the ethernet device. Routing is actually a very simple mechanism, but don’t worry if you find it slightly difficult to understand at first; everybody does.

You can display the contents of your routing table by using the route command without any options.

There are four commonly used routing mechanisms for unix network configurations. I’ll briefly discuss each in turn.

10.1 Static/Manual Routes.

Static routing, as its name implies, is ‘hard coded’ routing, that is, it will not change if your network suffers some failure, or if an alternate route becomes available. Static routes are often used in cases where you have a very simple network with no alternate routes available to a destination host, that is, there is only one possible network path to a destination host, or where you want to route a particular way to a host regardless of network changes.

In Linux there is a special use for manual routes, and that is for adding a route to a slip or plip host where you have used the ifconfig pointopoint parameter. If you have a slip/plip link, and have the pointopoint parameter specifying the address of the remote host, then you should add a static route to that address so that the ip routing software knows how to route datagrams to that address. The route command you would use for the slip/plip link via leased line example presented earlier would be:

     #/sbin/route add IPR.IPR.IPR.IPR

Where:

IPR.IPR.IPR.IPR represents the IP address of the remote end.

10.2 Default Route.

The default route mechanism is probably the most common and most useful to most end-user workstations and hosts on most networks. The default route is a special static route that matches every destination address, so that if there is no more specific route for a datagram to be sent to, then the default route will be used.

If you have a configuration where you have only a single ethernet interface, or a single slip interface device defined then you should point your default route via it. In the case of an ethernet interface, the Linux kernel knows where to send datagrams for any host on your network. It works this out using the network address and the network mask as discussed earlier. This means that the only datagrams the kernel won’t know how to properly route will be those for people not on your network. To make this work you would normally have your default route point to your router address, as it is your means of getting outside of your local network. If you are using a slip connection, then your slip server will be acting as your router, so your default route will be via your slip server.

To configure your default route, add the following to your rc.inet1 after all of your network device configurations:

     #
     # Add a default route.

     #
     /sbin/route add default gw RGA.RGA.RGA.RGA
     #

Where:

RGA.RGA.RGA.RGA represents your Router/Gateway Address.

10.3 Proxy ARP.

This method is ugly, hazard prone and should be used with extreme care, some of you will want to use it anyway.

Those with the greatest need for proxy arp will be those of you who are configuring your Linux machine as a slip dial-in server. For those of you who will be using PPP, the PPP daemon simplifies and automates this task, making it a lot safer to use.

Normally when a tcp/ip host on your ethernet network wants to talk to you, it knows your IP address, but doesn’t know what hardware (ethernet) address to send datagrams to. The ARP mechanism is there specifically to provide that mapping function between network address and hardware address. The ethernet protocol provides a special address that is recognised by all ethernet cards, this is called the broadcast address. ARP works by sending a specially formatted datagram containing the IP address of the host it wishes to discover the hardware address of, and transmits it to the ethernet broadcast address. Every host will receive this datagram and the host that is configured with the matching IP address will reply with its hardware address. The host that performed the arp will then know what hardware address to use for the desired IP address.

If you want to use your machine as a server for other machines, you must get your machine to answer ARP requests for their IP addresses on their behalf, as they will not be physically connected to the ethernet network. Lets say that you have been assigned a number of IP addresses on your local network that you will be offering to dial-in slip users. Lets say those addresses are: 128.253.154.120-124, and that you have an ethernet card with a hardware address of 00:00:C0:AD:37:1C. (You can find the hardware address of your ethernet card by using the ifconfig command with no options). To instruct your Linux server to answer arp requests by proxy for these addresses you would need to add the following commands to the end of your

rc.inet1 file:
     #
     # Proxy ARP for those dialin users who will be using this
     #           machine as a server:
     #
     /sbin/arp -s 128.263.154.120 00:00:C0:AD:37:1C pub
     /sbin/arp -s 128.263.154.121 00:00:C0:AD:37:1C pub
     /sbin/arp -s 128.263.154.122 00:00:C0:AD:37:1C pub
     /sbin/arp -s 128.263.154.123 00:00:C0:AD:37:1C pub
     /sbin/arp -s 128.263.154.124 00:00:C0:AD:37:1C pub

     #
     # End proxy arps.

The pub argument stands for ‘publish’. It is this argument that instructs your machine to answer requests for these addresses, even though they are not for your machine. When it answers it will supply the hardware address specified, which is of course its own hardware address.

Naturally you will need to ensure that you have routes configured in your linux server that point these addresses to the slip device on which they will be connecting.

If you are using PPP, you don’t need to worry about manually messing with the arp table, as the pppd will manage those entries for you if you use the proxyarp parameter, and as long as the IP addresses of the remote machine and the server machine are in the same network. You will need to supply the netmask of the network on the server’s pppd command line.

10.4 gated - the routing daemon.

gated could be used in place of proxy arp in some cases, and would certainly be much cleaner, but its primary use is if you want your linux machine to act as an intelligent ip router for your network. gated provides support for a number of routing protocols. Among these are RIP, BGP, EGP, HELLO, and OSPF. The most commonly used in small networks being rip. rip stands for ‘Routing Information Protocol’. If you run gated, configured for rip, your linux machine will periodically broadcast a copy of its routing table to your network in a special format. In this way, all of the other machines on your network will know what addresses are accessible via your machine.

gated can be used to replace proxy arp when all hosts on your network run either gated or routed. If you have a network where you use a mixture of manual and dynamic routes, you should mark any manual routes as passive to ensure that they aren’t destroyed by gated because it hasn’t received an update for them. The best way to add static routes if you are using gated is to add a static stanza to your /etc/gated.conf file. This is described below.

gated would normally be started from your rc.inet2 which is covered in the next section. You might already see a daemon called routed running. gated is superior to routed in that it is more flexible and more functional. So you should use gated and not routed.

10.4.1 Obtaining gated

Gated is available from:

sunsite.unc.edu

     /pub/Linux/system/Network/daemons/gated.linux.bin.tgz
                                      /gated.linux.man.tgz
                                      /gated.linux.tgz

gated.linux.tgz is the source, so you probably won’t need it unless you wish to recompile the binaries for some reason.

10.4.2 Installing gated

The gated binary distribution comprises three programs and two sample configuration files.

The programs are:

gated

the actual gated daemon.

gdc

the operational user interface for gated. gdc is for controlling the gated daemon, stopping and starting it, obtaining its status and the like.

ripquery

a diagnostic tool to query the known routes of a gateway using either a ‘rip query’ or a ‘rip poll’.

The configuration files are:

gated.conf

this is the actual configuration file for the gated daemon. It allows you to specify how gated will behave when it is running. You can enable and disable any of the routing protocols, and control the behaviour of those routing protocols running.

gated.version

a text file that describes the version number of the gated daemon

The gated binary distribution will not install the gated files in the correct place for you. Fortunately there aren’t very many, so its fairly simple to do.

To install the binaries try the following:

     # cd /tmp
     # gzip -dc .../gated.linux.bin.tgz | tar xvf -
     # install -m 500 bin/gated /usr/sbin
     # install -m 444 bin/gated.conf bin/gated.version /etc
     # install -m 555 bin/ripquery bin/gdc /sbin
     # rm -rf /tmp/bin

I keep the networking daemons in /usr/sbin, if yours are somewhere else then naturally you’ll have to change the target directory. The sample gated configuration file included configures gated to emulate the old routed daemon. It will probably work for you in most circumstances, and it looks like this:

     #
     #  This configuration emulates routed.  It runs RIP and only sends
     #  updates if there are more than one interfaces up and IP forwarding is
     #  enabled in the kernel.

     #
     #       NOTE that RIP *will not* run if UDP checksums are disabled in
     #       the kernel.
     #
     rip yes ;
     traceoptions all;
     #

If you have any static routes you wish to add, you can add them in a static stanza appended to your

/etc/gated.conf as follows:
     #
     static {
       37.0.0.0 mask 255.0.0.0 gateway 44.136.8.97 ;
       host 44.136.8.100 gateway 44.136.8.97 ;
     } ;
     #

The above example would create a static route to the Class A network 37.0.0.0 via gateway 44.136.8.97, and a static route to a host with address 44.136.8.100 via gateway 44.136.8.97. If you do this you do not need to add the routes using the route command, gated will add and manage the routes for you.

To install the man files, try the following:

     # cd /tmp
     # gzip -dc .../gated.linux.man.tgz | tar xvf -
     # install -m 444 man/*.8 /usr/man/man8
     # install -m 444 man/*.5 /usr/man/man5
     # rm -rf /tmp/man

The man files contain concise and detailed information on the configuration and use of gated. For information on configuring gated, refer to the gated-config man page.

11 Configuring the network daemons.

As mentioned earlier, there are other files that you will need to complete your network installation. These files concern higher level configurations of the network software. Each of the important ones are covered in the following sub-sections, but you will find there are others that you will have to configure as you become more familiar with the network suite.

11.1 /etc/rc.d/rc.inet2 (the second half of rc.net)

If you have been following this document you should at this stage have built an rc file to configure each of your network devices with the correct addresses, and set up whatever routing you will need for your particular network configuration. You will now need to actually start some of the higher level network software.

Now would be a really good time to read Olaf’s Network Administrators Guide, as it really should be considered the definitive document for this stage of the configuration process. It will help you decide what to include in this file, and more importantly perhaps, what not to include in this file. For the security conscious it is a fair statement to say that the more network services you have running, the more likely the chance of your system having a security hole: Run only what you need.

There are some very important daemons (system processes that run in the background) that you will need to know a little about. The man pages will tell you more, but they are:

11.1.1 inetd.

inetd is a program that sits in the background and manages internet connection requests and the like. It is smart enough that you don’t need to leave a whole bunch of servers running when there is nothing connected to them. When it sees an incoming request for a particular service, eg telnet, or ftp, it will check the /etc/services file, find what server program needs to be run to manage the request, start it, and hand the connection over to it. Imagine it as a master server for your internet servers. It also has a few simple standard services inbuilt. These are echo, discard and generate services used for various types of network testing. inetd doesn’t manage all servers and services that you might run, but it manages most of the usual ones. Normally services such as udp based services, or services that manage their own connection multiplexing such as World Wide Web servers or muds would be run independently of inetd. Generally the documentation accompanying such servers will tell you whether to use inetd or not.

11.1.2 syslogd.

syslogd is a daemon that handles all system logging. It accepts messages generated for it and will distribute them according to a set of rules contained in /etc/syslogd.conf. For example, certain types of messages you will want to send to the console, and also to a log file, where others you will want only to log to a file. syslogd allows you to specify what messages should go where.

11.2 A sample rc.inet2 file.

The following is a sample rc.inet2 file that Fred built. It starts a large number of servers, so you might want to trim it down to just those services that you actually want to run. To trim it down, simply delete or comment out the stanzas (if to fi) that you don’t need. All each stanza does is test that the relevant module is a file, that it exists, echoes a comment that you can see when you boot your machine, and then executes the commands with the arguments supplied to ensure that it runs happily in the background. For more detailed information on each of the deamons, check either the Network Administrators Guide or the relevant man pages.

     #! /bin/sh
     #
     # rc.inet2      This shell script boots up the entire INET system.
     #               Note, that when this script is used to also fire
     #               up any important remote NFS disks (like the /usr

11. Configuring the network daemons. 41

#               distribution), care must be taken to actually
#               have all the needed binaries online _now_...
#
# Version:      @(#)/etc/rc.d/rc.inet2  2.18    05/27/93
#
# Author:       Fred N. van Kempen, <waltje@uwalt.nl.mugnet.org>
#
# Constants.
NET="/usr/sbin"
IN_SERV="lpd"
LPSPOOL="/var/spool/lpd"
# At this point, we are ready to talk to The World...
echo -e "\nMounting remote file systems ..."
/bin/mount -t nfs -v            # This may be our /usr runtime!!!
echo -e "\nStarting Network daemons ..."
# Start the SYSLOG daemon.  This has to be the first server.
# This is a MUST HAVE, so leave it in.
echo -n "INET: "
if [ -f ${NET}/syslogd ]
then
        echo -n "syslogd "
        ${NET}/syslogd
fi
# Start the SUN RPC Portmapper.
if [ -f ${NET}/rpc.portmap ]
then
        echo -n "portmap "
        ${NET}/rpc.portmap
fi
# Start the INET SuperServer
# This is a MUST HAVE, so leave it in.
if [ -f ${NET}/inetd ]
then
        echo -n "inetd "
        ${NET}/inetd
else
        echo "no INETD found.  INET cancelled!"
        exit 1
fi
# Start the NAMED/BIND name server.
# NOTE: you probably don't need to run named.
#if [ ! -f ${NET}/named ]

#then
#        echo -n "named "
#        ${NET}/named
#fi
# Start the ROUTEd server.
# NOTE: routed is now obselete. You should now use gated.
#if [ -f ${NET}/routed ]
#then
#        echo -n "routed "
#        ${NET}/routed -q #-g -s
#fi
# Start the GATEd server.
if [ -f ${NET}/gated ]
then
        echo -n "gated "
        ${NET}/gated
fi
# Start the RWHO server.
if [ -f ${NET}/rwhod ]
then
        echo -n "rwhod "
        ${NET}/rwhod -t -s
fi
# Start the U-MAIL SMTP server.
if [ -f XXX/usr/lib/umail/umail ]
then
        echo -n "umail "
        /usr/lib/umail/umail -d7 -bd </dev/null >/dev/null 2>&1 &
fi
# Start the various INET servers.
for server in ${IN_SERV}
do
        if [ -f ${NET}/${server} ]
        then
                echo -n "${server} "
                ${NET}/${server}
        fi
done
# Start the various SUN RPC servers.
if [ -f ${NET}/rpc.portmap ]
then
        if [ -f ${NET}/rpc.ugidd ]

             then
                     echo -n "ugidd "
                     ${NET}/rpc.ugidd -d
             fi
             if [ -f ${NET}/rpc.mountd ]
             then
                     echo -n "mountd "
                     ${NET}/rpc.mountd
             fi
             if [ -f ${NET}/rpc.nfsd ]
             then
                     echo -n "nfsd "
                     ${NET}/rpc.nfsd
             fi
             # Fire up the PC-NFS daemon(s).
             if [ -f ${NET}/rpc.pcnfsd ]
             then
                     echo -n "pcnfsd "
                     ${NET}/rpc.pcnfsd ${LPSPOOL}
             fi
             if [ -f ${NET}/rpc.bwnfsd ]
             then
                     echo -n "bwnfsd "
                     ${NET}/rpc.bwnfsd ${LPSPOOL}
             fi
     fi
     echo network daemons started.
     # Done!

11.3 Other necessary network configuration files.

There are other network configuraiton files that you will need to configure if you want to have people connect to and use your machine as a host. If you have installed your linux from a distribution then you will probably already have copies of these files so just check them to make sure they look ok, and if not you can use the following samples.

11.3.1 A sample /etc/inetd.conf file.

Your /etc/rc.d/rc.inet2 file will have started inetd, syslogd and the various rpc servers for you. You will now need to configure the network daemons that will be managed by inetd. inetd uses a configuration file

called /etc/inetd.conf.

The following is an example of how a simple configuration might look:

     #

# The internal services.
#
# Authors:      Original taken from BSD UNIX 4.3/TAHOE.
#               Fred N. van Kempen, <waltje@uwalt.nl.mugnet.org>
#
echo    stream tcp nowait root  internal
echo    dgram  udp wait   root  internal
discard stream tcp nowait root  internal
discard dgram  udp wait   root  internal
daytime stream tcp nowait root  internal
daytime dgram  udp wait   root  internal
chargen stream tcp nowait root  internal
chargen dgram  udp wait   root  internal
#
# Standard services.
#
ftp     stream tcp nowait root  /usr/sbin/tcpd in.ftpd ftpd
telnet  stream tcp nowait root  /usr/sbin/tcpd in.telnetd
#
# Shell, login, exec and talk are BSD protocols.
#
shell   stream tcp nowait root  /usr/sbin/tcpd in.rshd
login   stream tcp nowait root  /usr/sbin/tcpd in.rlogind
exec    stream tcp nowait root  /usr/sbin/tcpd in.rexecd
talk    dgram  udp wait   root  /usr/sbin/tcpd in.talkd
ntalk   dgram  udp wait   root  /usr/sbin/tcpd in.talkd
#
# Status and Information services.
#
finger  stream tcp nowait root  /usr/sbin/tcpd in.fingerd
systat  stream tcp nowait guest /usr/sbin/tcpd /usr/bin/ps -auwwx
netstat stream tcp nowait guest /usr/sbin/tcpd /bin/netstat
#
# End of inetd.conf.

The inetd man page describes what each of the fields are, but put simply, each entry describes what program should be executed when an incoming connection is received on the socket listed as the first entry. Those entries which have incoming where the program name and arguments would be are those services that are provided internally by the inetd program.

The conversion between the service name in the first column, and the actual socket number it refers to is performed by the /etc/services file.

11.3.2 A sample /etc/services file.

The /etc/services file is a simple table of Internet service names and the socket number and protocol is uses. This table is used by a number of programs including inetd, telnet and tcpdump. It makes life a little easier by allowing us to refer to services by name rather than by number.

The following is a sample of what a simple /etc/services file might look like:

#
# /etc/services - database of service name, socket number
#                 and protocol.
#
# Original Author:
#     Fred N. van Kempen, <waltje@uwalt.nl.mugnet.org>
#
tcpmux     1/tcp
echo       7/tcp
echo       7/udp
discard    9/tcp   sink null
discard    9/udp   sink null
systat     11/tcp  users
daytime    13/tcp
daytime    13/udp
netstat    15/tcp
chargen    19/tcp  ttytst source
chargen    19/udp  ttytst source
ftp-data   20/tcp
ftp        21/tcp
telnet     23/tcp
smtp       25/tcp  mail
time       37/tcp  timserver
time       37/udp  timserver
name       42/udp  nameserver
whois      43/tcp  nicname    # usually to sri-nic
domain     53/tcp
domain     53/udp
finger     79/tcp
link       87/tcp  ttylink
hostnames  101/tcp hostname   # usually to sri-nic
sunrpc     111/tcp
sunrpc     111/tcp portmapper # RPC 4.0 portmapper TCP
sunrpc     111/udp
sunrpc     111/udp portmapper # RPC 4.0 portmapper UDP
auth       113/tcp authentication
nntp       119/tcp usenet     # Network News Transfer
ntp        123/tcp            # Network Time Protocol
ntp        123/udp            # Network Time Protocol
snmp       161/udp
snmp-trap  162/udp
exec       512/tcp            # BSD rexecd(8)
biff       512/udp comsat
login      513/tcp            # BSD rlogind(8)
who        513/udp whod       # BSD rwhod(8)

shell      514/tcp cmd        # BSD rshd(8)
syslog     514/udp            # BSD syslogd(8)
printer    515/tcp spooler    # BSD lpd(8)
talk       517/udp            # BSD talkd(8)
ntalk      518/udp            # SunOS talkd(8)
route      520/udp routed     # 521/udp too
timed      525/udp timeserver
mount      635/udp            # NFS Mount Service
pcnfs      640/udp            # PC-NFS DOS Authentication
bwnfs      650/udp            # BW-NFS DOS Authentication
listen     1025/tcp listener  # RFS remote_file_sharing
ingreslock 1524/tcp           # ingres lock server
nfs        2049/udp           # NFS File Service
irc        6667/tcp           # Internet Relay Chat
# End of services.

The telnet entry tells us that the telnet service uses socket number 23 and the tcp protocol. The domain entry tells us that the Domain Name Service uses socket number 52 and both tcp and udp protocols. You should have an appropriate /etc/services entry for each /etc/inetd.conf entry.

11.3.3 A sample /etc/protocols file.

The /etc/protocols file is a table of protocol name with its corresponding protocol number. Since the number of protocols in use is small this file is quite trivial.

#
# /etc/protocols - database of protocols.
#
# Original Author:
#   Fred N. van Kempen, <waltje@uwalt.nl.mugnet.org>
#
ip   0   IP   # internet protocol
icmp 1   ICMP # internet control message protocol
igmp 2   IGMP # internet group multicast protocol
ggp  3   GGP  # gateway-gateway protocol
tcp  6   TCP  # transmission control protocol
pup  12  PUP  # PARC universal packet protocol
udp  17  UDP  # user datagram protocol
idp  22  IDP
raw  255 RAW
#
# End of protocols.

11.4 Name Resolution.

Name Resolution is the process of converting a hostname in the familiar dotted notation (e.g. tsx-11.mit.edu) into an IP address which the network software understands. There are two principal means of achieving this in a typical installation, one simple, and one more complex.

11.4.1 /etc/hosts

/etc/hosts contains a list of ip addresses and the hostnames they map to. In this way, you can refer to other machines on the network by name, as well as their ip address. Using a nameserver (see section ‘named’) allows you to do the same name-¿ip address translation automatically. (Running named allows you to run your own nameserver on your linux machine). This file needs to contain at least an entry for 127.0.0.1 with the name localhost. If you’re not only using loopback, you need to add an entry for your ip address, with your full hostname (such as loomer.vpizza.com). You may also wish to include entries for your gateways and network addresses.

For example, if loomer.vpizza.com has the ip address 128.253.154.32, the /etc/hosts file would contain:

     # /etc/hosts
     # List of hostnames and their ip addresses
     127.0.0.1               localhost
     128.253.154.32          loomer.vpizza.com loomer
     # end of hosts

Once again you will need to edit this file to suit your own needs. If you’re only using loopback, the only line in /etc/hosts should be for 127.0.0.1, with both localhost and your hostname after it.

Note that in the second line, above, there are two names for 128.253.154.32: loomer.vpizza.com and just loomer. The first name is the full hostname of the system, called the ”Fully Qualified Domain Name”, and the second is an alias for it. The second allows you to type only rlogin loomer instead of having to type the entire hostname. You should ensure that you put the Fully Qualified Domain Name in the line before the alias name.

11.4.2 named - do I need thee ?

‘I dub thee ..’

named is the nameserver daemon for many unix-like operating systems. It allows your machine to serve the name lookup requests, not only for itself, but also for other machines on the network, that is, if another machine wants to find the address for `goober.norelco.com', and you have this machines address in your named database, then you can service the request and tell other machines what `goobers' address is.

Under older implementations of Linux tcp/ip, to create aliases for machine names, (even for your own machine), you had to run named on your Linux machine to do the hostname to IP address conversion. One problem with this is that named is comparitively difficult to set up properly, and maintain. To solve this problem, a program called hostcvt.build was made available on Linux systems to translate your /etc/hosts file into the many files that make up named database files. However even with this problem overcome, named still uses CPU overhead and causes network traffic.

The bottom line is this: You do not need to run named on your Linux system. The SLS instructions will probably tell you to run hostcvt.build to setup named. This is simply unnecessary unless you want to make your Linux system function as a nameserver for other machines, in which case you probably should learn some more about named anyway. When looking up hostnames, your linux machine will first check the /etc/hosts file, and then ask the nameserver out on the net.

The only reason you may want to run named would be if:

• You’re setting up a network of machines, and need a nameserver for one of them, and don’t have a nameserver out on the net somewhere.

• Your network administrators want you to run your Linux system as a nameserver for some reason.

• You have a slow slip connection, and want to run a small cache-only nameserver on your Linux machine so that you don’t have to go out on the serial line for every name lookup that occurs. If you’re only going to be connecting to a small number of hosts on the net, and you know what their addresses are, then you can put them in your hosts file and not need to query a nameserver at all. Generally namelookup isn’t that slow and should work fine over a slip link anyway.

• You want to run a nameserver for fun and excitement.

In general, you do NOT need to run named: this means that you can comment it out from your rc.inet2 file, and you don’t have to run hostcvt.build. If you want to alias machine names, for example, if you want to refer to loomer.vpizza.com as just loomer, then you can add as alias in /etc/hosts instead. There is no reason to run named unless you have a specific requirement to do so. If you have access to a nameserver, (and your network administrators will tell you its address), and most networks do, then don’t bother running named.

If you’re only using loopback, you can run named and set your nameserver address to 127.0.0.1, but since you are the only machine you can talk to, this would be quite bizzarre, as you’d never need to call it.

11.4.3 /etc/networks

The /etc/networks file lists the names and addresses of your own, and other, networks. It is used by the route command, and allows you to specify a network by name, should you so desire.

Every network you wish to add a route to using the route command should have an entry in the /etc/networks file, unless you also specify the -net argument in the route command line.

Its format is simliar to that of /etc/hosts file above, and an example file might look like:

     #
     # /etc/networks: list all networks that you wish to add route commands
     #                for in here
     #
     default         0.0.0.0         # default route    - recommended
     loopnet         127.0.0.0       # loopback network - recommended
     mynet           128.253.154.0   # Example network CHANGE to YOURS
     #
     # end of networks

11.4.4 /etc/host.conf

The system has some library functions called the resolver library. This file specifies how your system will lookup host names. It should contain at least the following two lines:

     order hosts,bind
     multi on

These two lines tell the resolve libraries to first check the /etc/hosts file, and then to ask the nameserver (if one is present). The multi entry allows you to have multiple IP addresses for a given machine name in

/etc/hosts.

This file comes from the implementation of the resolv+ bind library for Linux. You can find further documentation in the resolv+(8) man page if you have it. If you don’t, it can be obtained from:

sunsite.doc.ic.ac.uk

     /computing/comms/tcpip/nameserver/resolv+/resolv+2.1.1.tar.Z

This file contains the resolv+.8 man page for the resolver library.

11.4.5 /etc/resolv.conf

This file actually configures the system name resolver, and contains two types of entries: The addresses of your nameservers (if any), and the name of your domain, if you have one. If you’re running your own nameserver (i.e running named on your Linux machine), then the address of your nameserver is 127.0.0.1, the loopback address.

Your domain name is your fully qualified hostname (if you’re a registered machine on the Internet, for example), with the hostname component removed. That is, if your full hostname is loomer.vpizza.com, then your domain name is vpizza.com, without the hostname loomer.

For example, if you machine is goober.norelco.com, and has a nameserver at the address 128.253.154.5, then your /etc/resolv.conf file would look like:

     domain norelco.com
     nameserver 127.253.154.5

You can specify more than one nameserver. Each one must have a nameserver entry in the resolv.conf file.

Remember, if you’re running on loopback, you don’t need a nameserver.

11.4.6 Configuring your Hostname - /etc/HOSTNAME

After you have configured everything else, there is one small task that remains, you need to configure your own machine with a name. This is so that application programs like sendmail can know who you are to accept mail, and so that your machine can identify itself to other machines that it might be connected to.

There are two programs that are used to configure this sort of information, and they are commonly misused. They are hostname and domainname.

If you are using a release of net-tools earlier than 1.1.38 then you can include a command in your /etc/rc file that looks like this:

     /bin/hostname -S

and this will cause the hostname command to read a file called /etc/HOSTNAME which it expects will contain a ”Fully Qualified Domain Name”, that is, your machines hostname including the domainname. It will split the F.Q.D.N. into its DNS hostname and domainname components and set them appropriately for you.

For example, the machine above would have the file /etc/HOSTNAME:

     goober.norelco.com

If you are using the hostname that came with net-tools-1.1.38 or later, then you would add a command at the end of your /etc/rc.d/rc.inet1 file like:

     /bin/hostname goober.norelco.com

or if you have upgraded from a previous release, you could add:

     /bin/hostname -F /etc/HOSTNAME

and it would behave in the same way as for the earlier version.

The /bin/domainname command is for setting the N.I.S. domain name NOT the D.N.S. domain name. You do not need to set this unless you are running NIS, which is briefly described later.

11.5 Other files.

There are of course many other files in the /etc directory which you may need to dabble with later on. Instead of going into them here, I’m going to provide the bare minimum to get you on the net. More information is available in Olaf’s Network Administration Guide. It picks up where this HOWTO ends, and some more information will be provided in later versions of this document.

Once you have all of the files set up, and everthing in the right place, you should be able to reboot you new kernel, and net away to your hearts content. However I strongly suggest that you keep a bootable copy of your old kernel and possibly even a ‘recovery disk’, in case something goes wrong, so that you can get back in and fix it. You might try HJLu’s ‘single disk boot disk’, or ‘disk1’ from an SLS distribution.

12 Advanced Configurations.

The configurations above have described how a typical Linux workstation might be configured for normal end-user operation. Some of you will have other requirements which will require slightly more advanced configurations. What follows are examples of some the more common of these.

12.1 PPP - Point to Point Protocol.

The Point to Point Protocol is a modern and efficient protocol for conveying multiple protocols, tcp/ip for one, across serial links, that a lot of people use in place of slip. It offers enhanced functionality, error detection and security options. It corrects a number of deficiencies that are found in slip, and is suitable for both asynchronous links and synchronous links alike.

An important feature of PPP operation is dynamic address allocation, and this feature will almost certainly be exploited by your PPP server. This feature allows a PPP client, with a specially formatted frame, to request its address from the server. In this way configuration is somewhat less messy than with slip, since this ability to retrieve your address must occur outside of the protocol.

The authors of the Linux port are Michael Callahan, <callahan@maths.ox.ac.uk> and Al Longyear,

<longyear@netcom.com>. Most of this information has come from the documentation that accompanies the PPP software. The documentation is quite complete, and will tell you much more than I present here.

The Linux PPP code has come out of Alpha testing and is now available as a public release. The 1.0.0 Linux PPP code is based on Paul Mackerras’s free PPP for BSD-derivative operating systems. The 1.0.0 release is based on version 2.1.1 of the free PPP code.

The PPP code comes in two parts. The first is a kernel module which handles the assembly and disassembly of the frames, and the second is a set of protocols called LCP, IPCP, UPAP and CHAP, for negotiating link options, bringing the link into a functioning state and for authentication.

12.1.1 Why would I use PPP in place of SLIP ?

You would use PPP in place of SLIP for a few reasons. The most common are:

Your Internet Provider supports only PPP The most obvious reason you would use PPP in favour of SLIP is when your Internet Provider supports PPP and not SLIP. Ok, I said it was obvious.

You have a normally noisy serial line PPP provides a frame check sequence for each and every frame transmitted, SLIP does not. If you have a noisy serial line, and you are using SLIP, your error correction will be performed end to end, that is between your machine and the destination machine, whereas with PPP the error detection occurs locally, between your machine and the PPP server. This makes for faster recovery from errors.

You need to make use of some other feature PPP offers.

PPP provides a number of features that SLIP does not. You might for example want to carry not only IP, but also DECNET, or AppleTalk frames over your serial link. PPP will allow you to do this.

12.1.2 Where to obtain the PPP software.

The ppp software is available from:

sunsite.unc.edu

     /pub/Linux/system/Networking/serial/ppp-2.1.2b.tar.gz

This file contains the kernel source, and the pppd source and binary. Version 1.0.0 is meant for use with kernels 1.0.x and 1.1.x.

12.1.3 Installing the PPP software.

Installation of the PPP software is fairly straightforward.

The kernel driver. Some support for ppp has been built into the kernel for some time. Configuring the kernel is fairly easy, the following should work ok:

     # cd /usr/src
     # gzip -dc ppp-2.1.2b.tar.gz | tar xvf -

and if you are running a kernel prior to 1.1.14:

     # cp /usr/src/ppp-2.1.2b/linux/ppp.c /usr/src/linux/drivers/net
     # cp /usr/src/ppp-2.1.2b/pppd/ppp.h /usr/src/linux/include/linux

other wise do NOT copy these files as they will overwrite the ones in the kernel source.

If you are running a kernel version earlier than 1.1.13, or 1.0.x, then you will then need to uncomment the

CONFIG PPP line in /usr/src/linux/config.in.

If you are running a version of the kernel that is 1.1.3 or lower, then you will also need to uncomment out the macro definition of NET02D in the file /usr/src/linux/drivers/net/ppp.c by removing the /* characters.

You can then do:

     # make config   (remembering to answer yes to PPP support)
     # make dep
     # make

When you reboot with the new kernel you should see messages at boot time that look something like these:

     PPP: version 0.2.7 (4 channels) NEW_TTY_DRIVERS OPTIMIZE_FLAGS
     TCP compression code copyright 1989 Regents of the University of California
     PPP line discipline registered.

These indicate that the PPP support has in fact been compiled into your kernel.

Now, try looking at the contents of /proc/net/dev. It should look something like this:

       Inter-|   Receive                  |  Transmit
        face |packets errs drop fifo frame|packets errs drop fifo colls carrier
           lo:      0    0    0    0    0        0    0    0    0     0    0

         ppp0:      0    0    0    0    0        0    0    0    0     0    0
         ppp1:      0    0    0    0    0        0    0    0    0     0    0
         ppp2:      0    0    0    0    0        0    0    0    0     0    0
         ppp3:      0    0    0    0    0        0    0    0    0     0    0

This indicates that the kernel driver is installed correctly.

pppd If you want to recompile pppd, type make in the pppd subdirectory of the installation. There will be some warnings when compiling lcp.c, upap.c and chap.c but these are OK.

If you want to recompile chat, consult README.linux in the chat directory.

To install, type make install in the chat and pppd directories. This will put chat and pppd binaries in

/usr/sbin and the pppd.8 manual page in /usr/man/man8.

pppd needs to be run as root. You can either make it suid root or just use it when you are root. make install will try to install it suid root, so if you are root when you try to install it, it should work ok.

12.1.4 Configuring and using the PPP software.

Like slip, you can configure the PPP software as either a client or a server. The chat program performs a similar function to the dip program in that it is used to automate the dialling and login procedure to the remote machine, unlike dip though, it does not perform the ioctl to convert the serial line into a PPP line. This is performed by the pppd program. pppd can act as either the client or the server. When used as a client, it normally invokes the chat program to perform the connection and login, and then it takes over by performing the ioctl to change the line discipline to ppp, performs a number of steps in configuring your machine to talk to the remote machine and then steps out of the way to let you operate.

Please refer to the pppd and chat man pages for more information. Please also refer to the README file that comes with the ppp software, as its description of the operation of these utilities is much more complete than I have described here.

Configuring a PPP client by dial-up modem. This is perhaps what most of you will want to do, so it appears first. You would use this configuration when you have a network provider who supports ppp by dialup modem. When you want to establish your connection you simply have to invoke the pppd program with appropriate arguments.

The following example might look a little confusing at first, but it is easier to understand if you can see that all it is doing is taking a command line for the chat program as its first argument and then others for itself later.

     pppd connect 'chat -v "" ATDT5551212 CONNECT "" ogin: ppp word: password'\
           /dev/cua1 38400 debug crtscts modem defaultroute 192.1.1.17:

What this says is:

• Invoke the chat program with the command line:

          chat -v "" ATDT5551212 CONNECT "" ogin: ppp word: password

Which says: Dial 5551212, wait for the ‘CONNECT’ string, transmit a carriage return, wait for the string ‘ogin:’, transmit the string ‘ppp’, wait for the string ‘word:’, transmit the string ‘password’, and quit.

• Use serial device /dev/cua1

• Set its speed to 38400 bps.

• debug means log status messages to syslog

• crtscts means use hardware handshaking to the modem - recommended.

• modem means that pppd will attempt to hang up the call before and after making the call.

• defaultroute instructs pppd to add a routing entry that makes this the default route. In most cases this will be what you want.

• 192.1.1.17: says to set the ppp interfaces address to 192.1.1.17. This argument normally looks like x.x.x.x:y.y.y.y, where x.x.x.x is your ip address, and y.y.y.y is the ip address of the server. If you leave off the server’s address, pppd will ask for it, and x.x.x.x will be set to your machines ip address.

Please refer to the pppd and chat man pages for more information. Please also refer to the README file that comes with the ppp software, as its description of the above is much more complete than I have described here.

Configuring a PPP client via a leased line. Configuring a PPP client via a leased line is very simple. You will still use the pppd program, but since you won’t need to establish the modem link the arguments to the chat program can be much simpler.

The example I’m presenting here assumes that the ppp server doesn’t require any special login procedure. I do this because every login procedure will be different, and if you are simply running a local connection then it is possible that you might have it set up this way.

     pppd defaultroute noipdefault debug \
           kdebug 2 /dev/cua0 9600

This will open the serial device, generate the ioctl to change it into a pppdevice, set your default route via the ppp interface. The noipdefault argument instructs the pppd program to request the address to use for this device from the server. Debug messages will go to syslog. The kdebug 2 argument causes the debug messages to be set to level 2, this will give you slightly more information on what is going on. It will use /dev/cua0 at 9600 bps.

If your ppp server does require some sort of login procedure, you can easily use the chat program as in the example for the dialup server to perform that function for you.

Please refer to the pppd and chat man pages for more information. Please also refer to the README file that comes with the ppp software, as its description of the above is much more complete than I have described here.

Configuring a PPP server. Configuring a PPP server is similar to establishing a slip server. You can create a special ‘ppp’ account, which uses an executable script as its login shell. The /etc/passwd entry might look like:

     ppp:EncPasswd:102:50:PPP client login:/tmp:/etc/ppp/ppplogin

and the /etc/ppp/ppplogin script might look like:

     #!/bin/sh
     exec /usr/sbin/pppd passive :192.1.2.23

The address that you provide will be the address that the calling machine will be assigned.

Naturally, if you want multiple users to have simultaneous access you would have to create a number of startup scripts and individual accounts for each to use, as you can only put one ip address in each script.

12.1.5 Where to obtain more information on PPP, or report bugs.

Most discussion on PPP for Linux takes place on the PPP mailing list.

To join the Linux PPP channel on the mail list server, send mail to:

     linux-activists@niksula.hut.fi
     with the line:
     X-Mn-Admin: join PPP
     at the top of the message body (not the subject line).

Please remember that when you are reporting bugs or problems you should include as much information relevant to the problem as you can to assist those that will help you understand your problem.

You might also like to check out:

RFCS 1548, 1331, 1332, 1333, and 1334. These are the definitive documents for PPP.

W. Richard Stevens also describes PPP in his book ‘TCP/IP Illustrated Volume 1’, (Addison-Wessley, 1994, ISBN 0-201-63346-9).

12.2 Configuring Linux as a Slip Server.

If you have a machine that is perhaps network connected, that you’d like other people be able to dial into, and provide network services, then you will need to configure your machine as a server. If you want to use slip as the serial line protocol, then currently you have three options as to how to configure your Linux machine as a slip server. My preference would be to use the first presented, sliplogin, as it seems the easiest to configure and understand, but I will present a summary of each, so you make your mind.

12.2.1 Slip Server using sliplogin

sliplogin is a program that you can use in place of the normal login shell for slip users that converts the terminal line into a slip line. It allows you to configure your Linux machine as either a static address server, users get the same address everytime they call in, or a dynamic address server, where users get an address allocated for them which will not necessarily be the same as the last time they called.

The caller will login as per the standard login process, entering their username and password, but instead of being presented with a shell after their login, sliplogin is executed which searches its configuration file (/etc/slip.hosts) for an entry with a login name that matches that of the caller. If it locates one, it configures the line as an 8bit clean line, and uses an ioctl call to convert the line discipline to slip. When this process is complete, the last stage of configuration takes place, where sliplogin invokes a shell script which configures the slip interface with the relevant ip address, netmask and sets appropriate routing in place. This script is usually called /etc/slip.login, but in a similar manner to getty, if you have certain callers that require special initialisation, then you can create configuration scripts called /etc/slip.login.loginname that will be run instead of the default specifically for them.

There are either three or four files that you need to configure to get sliplogin working for you. I will detail how and where to get the software, and how each is configured in detail. The files are:

• /etc/passwd, for the dialin user accounts.

• /etc/slip.hosts, to contain the information unique to each dial-in user.

• /etc/slip.login, which manages the configuration of the routing that needs to be performed for the user.

• /etc/slip.tty, which is required only if you are configuring your server for dynamic address allocation and contains a table of addresses to allocate

• /etc/slip.logout, which contains commands to clean up after the user has hung up or logged out.

Where to get sliplogin sliplogin can be obtained from:

sunsite.unc.edu

     /pub/Linux/system/Network/serial/sliplogin-1.3.tar.gz

The tar file contains both source, precompiled binaries and a man page.

To ensure that only authorised users will be able to run sliplogin program, you should add an entry to your /etc/group file similar to the following:

      ..
     slip::13:radio,fred
      ..

When you install the sliplogin package, the Makefile will change the group ownership of the sliplogin program to slip, and this will mean that only users who belong to that group will be able to execute it. The example above will allow only users radio and fred to execute sliplogin.

To install the binaries into your /sbin directory, and the man page into section 8, do the following:

     # cd /usr/src
     # gzip -dc .../sliplogin-1.3.tar.gz | tar xvf -
     # cd src
     # make install

If you want to recompile the binaries before installation, add a make clean before the make install. If you want to install the binaries somewhere else, you will need to edit the Makefile install rule.

Please read the README files that come with the package for more information.

Configuring /etc/passwd for Slip hosts. Normally you wouldcreate some special logins for Slip callers in your /etc/passwd file. A convention commonly followed is to use the hostname of the calling host with a capital ‘S’ prefixing it. So, for example, if the calling host is called radio then you could create a /etc/passwd entry that looked like:

     Sradio:FvKurok73:1427:1:radio slip login:/tmp:/sbin/sliplogin

It doesn’t really matter what the account is called, so long as it is meaningful to you.

Note: the caller doesn’t need any special home directory, as they will not be presented with a shell from this machine, so /tmp is a good choice. Also note that sliplogin is used in place of the normal login shell.

Configuring /etc/slip.hosts The /etc/slip.hosts file is the file that sliplogin searches for entries matching the login name to obtain configuration details for this caller. It is this file where you specify the ip address and netmask that will be assigned to the caller, and configured for their use. Sample entries for two hosts, one a static configuration for host radio, and another, a dynamic configuration for user host albert might look like:

     #
     Sradio    44.136.8.99    44.136.8.100   0xffffff00   normal
     Salbert   44.136.8.99    DYNAMIC        0xffffff00   compressed
     #

The /etc/slip.hosts file entries are:

1. the login name of the caller.

2. ip address of the server machine, ie this machine.

3. ip address that the caller will be assigned. If this field is coded DYNAMIC then an ip address will be allocated based on the information contained in your /etc/slip.tty file discussed later. Note: you must be using at least version 1.3 of sliplogin for this to work.

4. the netmask assigned to the calling machine in hexadecimal notation eg 0xffffff00 for a Class C network mask.

5. optional parameters to enable/disable compression and other features.

Note: You can use either hostnames or IP addresses in dotted decimal notation for fields 2 and 3. If you use hostnames then those hosts must be resolvable, that is, your machine must be able to locate an ip address for those hostnames, otherwise the script will fail when it is called. You can test this by trying trying to telnet to the hostname, if you get the Trying nnn.nnn.nnn... message then your machine has been able to find an ip address for that name. If you get the message Unknown host, then it has not. If not, either use ip addresses in dotted decimal notation, or fix up your name resolver configuration (See section Name

Resolution).

The most commonly used optional paramaters for the opt1 and opt2 fields are:

normal

to enable normal uncompressed slip.

compressed

to enable van Jacobsen header compression (cslip)

Naturally these are mutually exclusive, you can use one or the other. For more information on the other options available, refer to the man pages.

Configuring the /etc/slip.login file. After sliplogin has searched the /etc/slip.hosts and found a matching entry, it will attempt to execute the /etc/slip.login file to actually configure the slip interface with its ip address and netmask.

The sample /etc/slip.login file supplied with the sliplogin package looks like this:

     #!/bin/sh -
     #
     #       @(#)slip.login  5.1 (Berkeley) 7/1/90
     #
     # generic login file for a slip line.  sliplogin invokes this with
     # the parameters:
     #     $1       $2        $3       $4         $5        $6    $7-n
     #   slipunit ttyspeed loginname local-addr remote-addr mask opt-args
     #
     /sbin/ifconfig $1 $4 pointopoint $5 mtu 1500 -trailers up
     /sbin/route add $5
     arp -s $5 <hw_addr> pub
     exit 0
     #

You will note that this script simply uses the ifconfig and route commands to configure the slip device with its ipaddress, remote ip address and netmask, and creates a route for the remote address via the slip device. Just the same as you would if you were using the slattach command.

Note also the use of Proxy ARP to ensure that other hosts on the same ethernet as the server machine will know how to reach the dial-in host. The <hw addr> field should be the hardware address of the ethernet card in the machine. If your server machine isn’t on an ethernet network then you can leave this line out completely.

Configuring the /etc/slip.logout file. When the call drops out, you want to ensure that the serial device is restored to its normal state so that future callers will be able to login correctly. This is achieved with the use of the /etc/slip.logout file. It is quite simple in format.

     #!/bin/sh -
     #
     #               slip.logout
     #
     /sbin/ifconfig $1 down
     /sbin/route del $5
     arp -d $5
     exit 0
     #

All it does is ‘down’ the interface and delete the manual route previously created. It also uses the arp command to delete any proxy arp put in place, again, you don’t need the arp command in the script if your server machine does not have an ethernet port.

Configuring the /etc/slip.tty file. If you are using dynamic ip address allocation (have any hosts configured with the DYNAMIC keyword in the /etc/slip.hosts file, then you must configure the /etc/slip.tty file to list what addresses are assigned to what port. You only need this file if you wish your server to dynamically allocate addresses to users.

The file is a table that lists the tty devices that will support dial-in slip connections and the ip address that should be assigned to users who call in on that port.

Its format is as follows:

     # slip.tty    tty -> IP address mappings for dynamic SLIP
     # format: /dev/tty?? xxx.xxx.xxx.xxx
     #
     /dev/ttyS0      192.168.0.100
     /dev/ttyS1      192.168.0.101
     #

What this table says is that callers that dial in on port /dev/ttyS0 who have their remote address field in the /etc/slip.hosts file set to DYNAMIC will be assigned an address of 192.168.0.100.

In this way you need only allocate one address per port for all users who do not require an dedicated address for themselves. This helps you keep the number of addresses you need down to a minimum to avoid wastage.

12.2.2 Slip Server using dip.

Let me start by saying that some of the information below came from the dip man pages, where how to run Linux as a slip server is briefly documented. Please also beware that the following has been based on the dip337j-uri.tgz package and probably will not apply to other versions of dip.

dip has an input mode of operation, where it automatically locates an entry for the user who invoked it and configures the serial line as a slip link according to information it finds in the /etc/diphosts file. This input mode of operation is activated by invoking dip as diplogin. This therefore is how you use dip as a slip server, by creating special accounts where diplogin is used as the login shell.

The first thing you will need to do is to make a symbolic link as follows:

     # ln -sf /usr/sbin/dip /usr/sbin/diplogin

You then need to add entries to both your /etc/passwd and your /etc/diphosts files. The entries you need to make are formatted as follows:

To configure Linux as a slip server with dip, you need to create some special slip accounts for users, where dip (in input mode) is used as the login shell. A suggested convention is that of having all slip accounts begin with a capital ‘S’, eg ‘Sfredm’.

A sample /etc/passwd entry for a slip user looks like:

     Sfredm:ij/SMxiTlGVCo:1004:10:Fred:/tmp:/usr/sbin/diplogin
     ^^^^^^^^^^^^^^
     |          |         |   |    |    |    \__diplogin as login shell
     |          |         |   |    |    \_______Home directory
     |          |         |   |    \____________User Full Name
     |          |         |   \_________________User Group ID
     |          |         \_____________________User ID
     |          \_______________________________Encrypted User Password
     \__________________________________________Slip User Login Name

After the user logs in, the login(1) program, if it finds and verifies the user ok, will execute the diplogin command. dip, when invoked as diplogin knows that it should automatically assume that it is being used a login shell. When it is started as diplogin the first thing it does is use the getuid() function call to get the userid of whoever has invoked it. It then searches the /etc/diphosts file for the first entry that matches either the userid or the name of the tty device that the call has come in on, and configures itself appropriately. By judicious decision as to whether to give a user an entry in the diphosts file, or whether to let the user be given the default configuration you can build your server in such a way that you can have a mix of static and dynamically assigned address users.

dip will automatically add a ‘Proxy-ARP’ entry if invoked in input mode, so you do not need to worry about manually adding such entries.

Configuring /etc/diphosts /etc/diphosts is used by dip to lookup preset configurations for remote hosts. These remote hosts might be users dialing into your linux machine, or they might be for machines that you dial into with your linux machine.

The general format for /etc/diphosts is as follows:

      ..
     Suwalt::145.71.34.1:145.71.34.2:255.255.255.0:SLIP uwalt:CSLIP,1006
     ttyS1::145.71.34.3:145.71.34.2:255.255.255.0:Dynamic ttyS1:CSLIP,296
      ..

The fields are:

1. login name: as returned by getpwuid(getuid()) or tty name.

2. unused: compat. with passwd

3. Remote Address: IP address of the calling host, either numeric or by name

4. Local Address: IP address of this machine, again numeric or by name

5. Netmask: in dotted decimal notation

6. Comment field: put whatever you want here.

7. protocol: Slip, CSlip etc.

8. MTU: decimal number

An example /etc/net/diphosts entry for a remote slip user might be:

     Sfredm::145.71.34.1:145.71.34.2:255.255.255.0:SLIP uwalt:SLIP,296

which specifies a slip link with remote address of 145.71.34.1, and MTU of 296, or:

     Sfredm::145.71.34.1:145.71.34.2:255.255.255.0:SLIP uwalt:CSLIP,1006

which specifies a cslip-capable link with remote address 145.71.34.1, and MTU of 1006.

Therefore, all users who you wish to be allowed a statically allocated dial-up IP access should have an entry in the /etc/diphosts and if you want users who call a particular port to have their details dynamically allocated you must have an entry for the tty device and do not configure a user based entry. You should remember to configure at least one entry for each tty device that your dialup users use to ensure that a suitable configuration is available for them regardless of which modem they call in on.

When a user logs in, they will receive a normal login and password prompt, at which they should enter their slip-login userid and password. If they check out ok, then the user will see no special messages, they should just change into slip mode at their end, and then they should be able to connect ok, and be configured with the parameters from the diphosts file.

12.2.3 slip servers using the dslip package.

Matt Dillon <dillon@apollo.west.oic.com> has written a package that does not only dial-in but also dial-out slip. Matt’s package is a combination of small programs and scripts that manage your connections for you. You will need to have tcsh installed as at least one of the scripts requires it. Matt supplies a binary copy of the expect utility as it too is needed by one of the scripts. You will most likely need some experience with expect to get this package working to your liking, but don’t let that put you off.

Matt has written a good set of installation instructions in the README file, so I won’t bother repeating them.

You can get the dslip package from its home site at:

apollo.west.oic.com

     /pub/linux/dillon_src/dslip203.tgz

or from:

sunsite.unc.edu

     /pub/Linux/system/Network/serial/dslip203.tgz

Read the README file, and create the /etc/passwd and /etc/group entries before doing a make install.

12.3 Using the Automounter Daemon - AMD.

This section has been supplied by Mitch DSouza, and I’ve included it with minimal editing, as he supplied it. Thanks Mitch.

12.3.1 What is an automounter, and why would I use one ?

An automounter provides a convenient means of mounting filesystems on demand, i.e. when requried. This will reduce both the server and the client load, and provides a great deal of flexibility even with non-NFS mounts. It also offers a redundancy mechanism whereby a mount point will automatically switch to a secondary server should a primary one be unavailable. A rather useful mount called the union mount gives the automounter the ability to merge the contents of multiple directories into a single directory. The documentation msut be read thoroughly to make full use of its extensive capabilities.

A few important points must be remembered - (in no particular order):

• amd maps are not compatible with Sun maps, which in turn are not compatible with HP maps ad infinitum. The point here however is that amd is freely available and compatible with all the systems mentioned above and more, thus giving you the ability to share maps if amd is installed throughout your network. Mitch uses it with a mixture of Linux/Dec/NeXt/Sun machines.

• Sun automount maps can be converted to amd style maps by using the perl script in the contrib

     directory - automount2amd.pl.

• You must have the portmapper running before starting amd.

• UFS mounts do not timeout.

• UFS mounts, in the case of Linux only, have been extended to deal with all varieties of native filesystems (i.e. minix, ext, ext2, xiafs ...) with the default being minix. This undocumented feature is accessed in the opts option like:

           ..., opts:=type=msdos,conv=auto

• Do not mount over existing directories unless you use a direct automount option, otherwise it is like mounting your disk on /home when some user directory is /home/fred.

• Always turn on full logging with the ‘-x all’ option to amd if you have any troubles. Check also what the command:

          % amq -ms

reports, as it will indicate problems as they occur.

• GNU getopt() is too clever for its own good sometimes. You should always use ‘–’ before the non-options e.g.

          # /etc/amd -x all -l syslog -a /amd -- /net /etc/amd.net

12.3.2 Where to get AMD, the automounter daemon.

amd can be obtained from:

sunsite.unc.edu

     /pub/Linux/system/Misc/mount/amd920824upl67.tar.gz

This contains ready-to-run binaries, full sources and documentation in texinfo format.

12.3.3 An example AMD configuration.

You do not configure the automounter from the /etc/fstab file, which you will already be using to contain information about your fileystems, instead it is command line driven.

To mount two nfs filesystems using your /etc/fstab file you would use two entries that looked like:

     server-1:/export/disk  /nfs/server-1  nfs  defaults
     server-2:/export/disk  /nfs/server-2  nfs  defaults

i.e. you were nfs mounting server-1 and server-2 on your linux disk on the /nfs/server-1 and

/nfs/server-2 directories.

After commenting out, or deleting the above lines from your /etc/fstab file, you could amd to perform the same task with the following syntax:

   /etc/amd -x all -l syslog -a /amd -- /nfs /etc/amd.server
   |      | |    | |       | |     |  | |  | |             |
   |      | |    | |       | |     |  | |  | |             |
   `------' `----' `-------' `-----' -' `--' `-------------'
    |        |      |         |      |   |    |
    (1)      (2)    (3)       (4)    (5) (6)  (7)

Where:

1. The full amd binary path (obviously optional) depending on your $PATH setting, so just ‘amd’ may be specified here.

2. ‘-x all’ means turn full logging on. Read the documentation for the other logging levels

3. ‘-l syslog’ means log the message via the syslog daemon. This could mean put it to a file, dump it, or pass it, to an unused tty console. This (syslog) can be changed to the name of a file, i.e. ‘-l foo’ will record to a file called foo.

4. ‘-a /amd’ means use the /amd directory as a temporary place for automount points. This directory is created automatically by amd and should be removed before starting amd in your /etc/rc scripts.

5. ‘–’ means tell getopt() to stop attempting to parse the rest of the command line for options. This is especially useful when specifying the ‘type:=’ options on the command line, otherwise getopt() tries to decode it incorrectly.

6. ‘/nfs’ is the real nfs mount point. Again this is automatically created and should not generally contain subdirectories unless the ‘type:=direct’ option is used.

7. The amd map (i.e. a file) named ‘amd.server’ contains the lines:

       # /etc/amd.server
       /defaults    opts:=rw;type:=nfs
       server-1     rhost:=server-1;rfs:=/export/disk
       server-2     rhost:=server-2;rfs:=/export/disk

Once started and successfully running, you can query the status of the mounts with the command:

   % amq -ms

Now if you say:

   % ls /nfs

you should see no files. However the command:

   % ls /nfs/server-1

will mount the host ‘server-1’ automatically. voila! amd is running. After the default timeout has expired, this will automatically be unmounted. Your /etc/password file could contain entries like:

      ...
     linus:EncPass:10:0:God:/nfs/server-1/home/linus:/bin/sh
     mitch:EncPass:20:10:Mitch DSouza:/nfs/server-1/home/mitch:/bin/tcsh
     matt:EncPass:20:10:Matt Welsh:/nfs/server-1/home/matt:/bin/csh

which would mean that when Linus, Matt, or Mitch are logged in, their home directory will be remotely mounted from the appropriate server, and umounted when they log out.

12.4 Using Linux as a router

Linux will function just fine as a router. You should run a routing daemon such as gated, or if you have simple routing requirements use hard coded routes. If you are using a late version kernel (1.1.*) then you should ensure that you have answered ‘y’ to:

     IP forwarding/gatewaying (CONFIG_IP_FORWARD) [y] y

when building your kernel.

Olaf Kirch’s Network Administrators Guide discusses network design and routing issues, and you should read it for more information. A reference to it is in the ”Related Documentation” section of this document.

12.5 NIS - Sun Network Information System.

There is now an NIS-HOWTO (http://sunsite.unc.edu/mdw/HOWTO/NIS-HOWTO.html) which you should

read if you are interested in using NIS. It details how to obtain, install and configure the NIS system for Linux.

13 Experimental and Developmental modules.

There are a number of people developing new features and modules for the Linux networking code. Some of these are in quite an advanced state (read working), and it is these that I intend to include in this section until they are standard release code, when they will be moved forward.

13.1 AX.25 - A protocol used by Amateur Radio Operators.

The AX.25 protocol is used by Amateur Radio Operators worldwide. It offers both connected and connectionless modes of operation, and is used either by itself for point-point links, or to carry other protocols such as tcp/ip and netrom.

It is similar to X.25 level 2 in structure, with some extensions to make it more useful in the amateur radio environment.

Alan Cox has developed some kernel based AX.25 software support for Linux and these are available in ALPHA form for you to try. Alan’s code supports both KISS based TNC’s (Terminal Node Controllers), and the Z8530 SCC driver.

The User programs contain a P.M.S. (Personal Message System), a beacon facility, a line mode connect program, and ‘listen’ an example of how to capture all AX.25 frames at RAW interface level.

Be sure to read /usr/local/ax25/README as it contains more complete information regarding this software.

13.1.1 Where to obtain the AX.25 software.

The AX.25 software is available from:

sunacm.swan.ac.uk

     /pub/misc/Linux/Radio/*

You will find a number of directories, each containing different versions of the code. Since it is closely linked with the kernel code, you will need to ensure that you choose the version appropriate for the kernel version you are running. The following table shows the mapping between the two:

     AX25007                 Prehistoric
     AX25010                 Obsolete
     AX25012                 for release 1.0.* kernels and higher
     AX25016                 for release 1.1.5 kernels
     AX25017                 for release 1.1.6 kernels
     AX25018
     AX25021
     AX25022                 for release 1.1.28 kernels

In each directory you will find at least two files, one called something like krnl022.tgz, and the other called something like user022.tgz. These are the kernel software, and the user programs respectively.

13.1.2 Installing the AX.25 software.

The software comes in two parts, the kernel drivers, and the user programs.

The kernel drivers. To install the kernel drivers, do the following:

     # cd /usr/src
     # gzip -dc krnl022.tgz | tar xvf -

you will need to uncomment the CONFIG AX25 define in the /usr/src/linux/config.in file.

You should then:

     # cd /usr/src/linux
     # make config
     # make dep;make

Be sure to answer ‘yes’ when you are asked if you should include the AX.25 support in the make config step. You will also need to answer ‘yes’ to inluding SLIP if you want the AX.25 code to support a KISS TNC.

The user programs. To install the user programs you should try:

     # cd /
     # gzip -dc user022.tgz | tar xvvof -

You should then:

     # cd /usr/local/ax25/src
     # make install

13.1.3 Configuring and using the AX.25 software.

Configuring an AX.25 port is very similar to configuring a slip device. The AX.25 software has been designed to work with a TNC in kiss mode. You will need to have the TNC preconfigured and connected.

You use the axattach program in much the same way as you would use the slattach program. For example:

     # /usr/local/ax25/bin/axattach -s 4800 /dev/cua1 VK2KTJ &

would configure your /dev/cua1 serial device to be a kiss interface at 4800 bps, with the hardware address

VK2KTJ.

You would then use the ifconfig program to configure the ip address and netmask as for an ethernet device:

     # /sbin/ifconfig sl0 44.136.8.5
     # /sbin/ifconfig sl0 netmask 255.255.255.0
     # /sbin/ifconfig sl0 broadcast 44.136.8.255
     # /sbin/ifconfig sl0 arp mtu 257 up

To test it out, try the following:

     /usr/local/ax25/bin/call VK2DAY via VK2RVT

The call program is a linemode terminal program for making ax.25 calls. It recognises lines that start with ‘ ’ as command lines. The ‘ .’ command will close the connection.

You also need to configure some items such as the window to use. This necessitates editing only one file. Edit the /usr/local/ax25/etc/ports file. This is an ascii file containing one line for each AX.25 port. You must have the entries in this file in the same order as you configure your AX.25 interfaces.

The format is as follows:

     callsign baudrate window frequency

At this stage not much of this information is used, it will be picked up and used in later developments.

I haven’t had a chance to try this code out yet. Please refer to the man pages in /usr/local/ax25/man and the README file in /usr/local/ax25 for more information.

13.2 Z8530 SCC driver.

The Zilog Z8530 SCC provides Synchronous/Asynchronous, HDLC, NRZI encoding and other capabilities. There are a number of peripheral cards that use the Z850 as the basis of their design. A driver has been

written by Joerg Reuter, <dl1bke@melaten.ihf.rwth-aachen.de>, and is available on:

ftp.ucsd.edu

     /hamradio/packet/tcpip/incoming/sccdrv-1.4a.dl1bke.tar.gz

Please read the README file that accompanies the driver for more details.

13.3 Ottawa PI/PI2 card driver.

The Ottawa PI card is a Z8530 SCC based card for IBM PC type machines that is in common usage by Amateur Radio operators worldwide. While it is most commonly used by Amateur Radio Operators, it could be pressed into service in other fields where it is desirable to have the features of a Z8530. It supports a high speed half duplex (single DMA channel) port, and a low speed (<19.2kbps interrupt driven) full duplex port. The PI2 is a new version of the card that supports an on board radio modem, and improved hardware design.

A driver for this card has been written by David Perry, <dp@hydra.carleton.edu>, and is available from:

hydra.carleton.ca

     /pub/hamradio/packet/tcpip/linux/pi2-0.5ALPHA.tgz

Please read the README file that accompanes the driver for more details.

13.4 snmp agent.

There is an experimental snmp agent for linux, ported by Erik Schoenfelder,

<schoenfr@ibr.cs.tu-bs.de>.

It is available from:

ftp.ibr.cs.tu-bs.de

     /pub/local/cmu-snmp2.1.2l2.tar.gz

Please read the file called cmu-snmp2.1.2l2.README, as it contains information that you will need to know about the package.

This package provides a nearly complete MIB-II variable set. At this stage though, you can only read variables, not set them.

nstat.tar.gz contains a formatter of the output from /proc/net/snmp called nstat.

You will need to be running either a new version kernel, or apply patches to your kernel source. Details are in the README file.

13.5 Experimental ARCNet driver

There have been a number of people looking for a driver for ARCNet cards. ARCNet cards provide only rates of about 2Mbps, but are capable of being supported via longer length cables than for 10base2 (thinnet) type lans. ARCNet cards are also likely to be fairly cheap, as many businesses are gradually replacing their ARCNet networks with other lan types.

Avery Pennarun <apenwarr@tourism.807-city.on.ca> has produced an experimental ARCNet driver for

Linux. It is ALPHA which means you should try it knowing that it probably still contains many errors and might even cause you other problems like kernel hangs.

The source code and kernel patch for the driver are available at:

sunsite.unc.edu

     /pub/Linux/system/Network/drivers/arcnet-0.22.tar.gz

Note: For ease of patching you will require kernel version 1.1.51 as this is what the patch was made against, though it might also work either 1.1.45 and up.

There are some known bugs, you can obtain details of these by reading the top of the arcnet.c file.

Avery is now at the stage where he needs people to try the driver to discover any bugs or problems that he hasn’t been able to find, and to guage how well it works in other environments. Avery will happily accept any reports via the NET channel, or by email to home at either <apenwarr@tourism.807-city.on.ca¿, or <Avery.Pennarun@NorLinK.Com¿. Avery has problems with News, so bug reports via any of the Linux newsgroups might not make it to him. Avery wants to know not only if you have problems, but also if you have success.

A mail list now exists to discuss ARCNet driver development. To subscribe, send a message with the following in the body of the text:

     subscribe linux-arcnet YOUR REAL NAME

13.6 Experimental Token Ring driver

An  experimental  Token  Ring  driver  is  being  developed  by  Peter

De Schrijver <stud11@cc4.kuleuven.ac.be>. His latest version, at the time of writing was available at: linux3.cc.kuleuven.ac.be

     /pub/Linux/TokenRing.patch-1.1.64.gz

ftp.cs.kuleuven.ac.be

     /pub/unix/linux/TokenRing.patch-1.1.64.gz

The patch file is against kernel version 1.1.64, as reflected in the patch filename.

Most boards based on IBM’s TROPIC chipset should work now. The following boards are known to be working with the driver :

• IBM Token Ring Adapter II

• IBM Token Ring 16/4 Adapter

• IBM Token Ring Adapter/A

• IBM Token Ring 16/4 Adapter/A

• HyperRing Classic 16/4

Boards which use the TI chipset or busmastering DMA won’t work with the current driver. However someone is working on a driver for the IBM busmaster adapters.

13.7 V.35 interface board

V.35 is a C.C.I.T.T. standard interface that provides a high speed balanced serial interface suitable for speeds up to about 2 Mbps. The use of differential pair balanced transmission allows the V.35 interface to support longer cables than can the more familiar V.24/RS232C type interface and higher data rates.

Pete Kruckenberg <kruckenb@sal.cs.utah.edu> located a company that supplies V.35 interface hardware for ISA bus machines. The company is also developing a Linux driver for this card that is nearing Beta testing stage. This would allow you to directly connect your Linux machine to a 48/56kbps synchronous leased line. The card supports multiple protocols and allows for interface speeds of up to 12 Mbps.

More information is available from:

ftp.std.com

     pub/sdl/n2
or you can email Dale Dhillon <sdl@world.std.com>

13.8 IPX bridge program

Vinod G Kulkarni <vinod@cse.iitb.ernet.in> has cowritten some software for linux that will allow it to act as an IPX bridge. The software is available from: sunsite.unc.edu

     /pub/Linux/Systems/Network/router/ipxbridge.tar.gz

13.9 IPX RIP and SAP support.

Alex Liu <labrat@unitrx.com has written support for the Novell RIP and SAP protocols to allow your linux machine to act as a Novell router. This software is alpha and includes a kernel patch. Be warned that you should take the usual precautions when testing this software. You can obtain the software from: sunsite.unc.edu

     /pub/Linux/Incoming/ipxripd-002.tar.gz (until it is moved)
     /pub/Linux//system/Network/router/ipxripd-002.tar.gz

A README file is included, and you should read this for installation and configuration details.

13.10 Demand Dial SLIP/PPP package

Eric Schenk <schenk@cs.toronto.edu> has written a demand dial daemon that will work with either SLIP or PPP. It relies on you having a slip device configured which the daemon connects to via a pty. When your slip connection is not active all datagrams for non local hosts will be routed to this device, and the daemon will detect them, when it receives a datagram it executes a script to activate your network link, and then reroutes datagrams to that link. The softwware is available at: sunsite.unc.edu

     /pub/Linux/system/Network/serial/diald-0.4.tar.gz

Note: You must configure your kernel so that it includes the slip driver, even if you only want to run PPP. The included documentation describes how to install and configure the software.

14 Diagnostic tools - How do I find out what is wrong?

In this section I’ll briefly describe some of the commonly used diagnostic tools that are available for your Linux network, and how you might use them to identify the cause of your network problems, or to teach yourself a bit more about how tcp/ip networking works. I’ll gloss over some of the detail of how the tools work because this document is not an appropriate forum for describing that sort of detail, but I hope I’ll have presented enough information that you’ll have an understanding of how to use the tool, and to better understand the relevant man page or other documentation.

14.1 ping - are you there?

The ping tool is located in the NetKit-B distribution as detailed above in the ‘Network Applications’ section. ping, as the name implies, allows you to transmit a datagram at another host that it will reflect back at you if it is alive and working ok and the network in between is also ok. In its simplest form you would simply say:

     # ping gw
     PING gw.vk2ktj.ampr.org (44.136.8.97): 56 data bytes
     64 bytes from 44.136.8.97: icmp_seq=0 ttl=254 time=35.9 ms
     64 bytes from 44.136.8.97: icmp_seq=1 ttl=254 time=22.1 ms
     64 bytes from 44.136.8.97: icmp_seq=2 ttl=254 time=26.0 ms
     ^C
     --- gw.vk2ktj.ampr.org ping statistics ---
     3 packets transmitted, 3 packets received, 0% packet loss
     round-trip min/avg/max = 22.1/28.0/35.9 ms
     #

What ping has done is resolved the hostname to an address, and using the icmp protocol has transmitted an icmp echo request datagram to the remote host periodically. For each echo request that the remote host receives it will formulate an icmp echo reply datagram which it will transmit back to you. Each line beginning with ‘64 bytes from ...’ represents an echo reply received in response to an echo request. Each line tells you the address of the host that sent you the reply, the sequence number to which the reply was for, the time to live field and the total round trip time that was taken. The round trip time is the time between when the echo request datagram is transmitted, and the corresponding echo reply is received. This can be used as a measure of how fast or slow the network connection between the two machines is.

The last two lines tell you how many datagrams were transmitted, how many valid responses were received and what percentage of the datagrams were lost. The percentage lost figure is a measure of how good or error free the network connection is. High percentage lost figures indicate such problems as a high error rate on a link somewhere between the hosts, exhausted capacity on a router or link somewhere, or high collision rate on an ethernet lan. You can use ping to identify where this problem might be by running ping sessions to each of the routed points that make up the network path. When you find that you can ping somewhere without any datagram loss, but pinging anywhere past there causes you packet loss, you can deduce that the problem lies somewhere between those two points.

14.2 traceroute - How do I get there?

The traceroute tool is found in the NetKit-A distribution detailed earlier. traceroute is primarily used for testing and displaying the path that your network connection would take to a destination host. traceroute also uses the icmp protocol, but it uses a clever trick to get each point along the path to send it back a reply as it creeps its way along. Its trick is to manually manipulate the time to live field of the datagrams it transmits. The time to live field is a mechanism that ensures that rogue datagrams do not get caught in a routing loop. Each time a datagram passes through a router it decrements the time to live field by one. If the time to live reaches zero then that router or host sends an icmp time to live expired message back to the host who transmitted the datagram to let it know the datagram has expired. traceroute uses this mechanism by sending a series of udp datagrams with the time to live beginning set at one, and incrementing each step it takes. By recording the addresses from the icmp time to live expired replies it receives in response to the datagrams dying it can determine the path taken to get to the destination. An example of its use would look something like:

     # traceroute minnie.vk1xwt.ampr.org
     traceroute to minnie.vk1xwt (44.136.7.129), 30 hops max, 40 byte packets
      1  gw (44.136.8.97)  51.618 ms  30.431 ms  34.396 ms
      2  gw.uts (44.136.8.68) 2017.322 ms  2060.121 ms 1997.793 ms
      3  minnie.vk1xwt (44.136.7.129) 2205.335 ms  2319.728 ms  2279.643 ms
     #

The first column tells us how many hops away (what the ttl value was), the second column is the hostname and address that responded if it could be resolved or just its address if it could not. The third, fourth and fifth columns are the round trip time for three consecutive datagrams to that point. This tells us that the first hop in the network route is via gw.vk2ktj, and the three figures following are the round trip times to that router. The next hop was via gw.uts.ampr.org, and minnie.vk1xwt.ampr.org is one hop further away. You can deduce information about the network route by looking at the difference in times between each step in the route. You can see that the round trip times to gw are fairly fast, it is an ethernet connected host. gw.uts is substantially slower to get to than gw, it is across a low speed radio link, so you have the ethernet time plus the radio link time added together. minnie.vk1xwt is only slightly slower than gw.uts, they are connected via a high speed network.

If you perform a traceroute and you see the string !N appear after the time figure, this indicates that your traceroute program received a network unreachable response. This message tells you that the host or router who sent you the message did not know how to route to the destination address. This normally indicates that there is a network link down somewhere. The last address listed is as far as you get before you find the faulty link.

Similarly if you see the string !H this indicates that a host unreachable message has been received. This might suggest that you got as far as the ethernet that the remote host is connected to, but the host itself is not responding or is faulty.

14.3 tcpdump - capturing and displaying network activity.

Adam Caldwell <acaldwel@103mort2.cs.ohiou.edu¿ has ported the tcpdump utlility to linux. tcpdump allows you to take traces of network activity by intercepting the datagrams on their way in and out of your machine. This is useful for diagnosing difficult to identify network problems.

Both binary and sources are available, and version 3.0 has been tested on kernel versions 0.99.15, 1.0.8 and 1.1.28.

You  can  find  the  source  and  binaries  at:  103mor2.cs.ohiou.edu

(ftp://103mort2.cs.ohiou.edu/linux/tcpdump-3.0-linux-src.tar.gz) or from: sunsite.unc.edu
(ftp://sunsite.unc.edu/pub/Linux/system/Network/tcpdump-3.0-linux-src.tar.gz)

tcpdump decodes each of the datagrams that it intercepts and displays them in a slightly cryptic looking format in text. You would use tcpdump if you were trying to diagnose a problem like protocol errors, or strange disconnections, as it allows you to actually see what has happened on the network. To properly use tcpdump you would need some understanding of the protocols and how they work, but it is useful for simpler duties such as ensuring that datagrams are actually leaving your machine on the correct port if you are trying to diagnose routing problems and for seeing if you are receiving datagrams from remote destinations.

A sample of tcpdump output looks like this:

     # tcpdump -i eth0
     tcpdump: listening on eth0
     13:51:36.168219 arp who-has gw.vk2ktj.ampr.org tell albert.vk2ktj.ampr.org
     13:51:36.193830 arp reply gw.vk2ktj.ampr.org is-at 2:60:8c:9c:ec:d4
     13:51:37.373561 albert.vk2ktj.ampr.org > gw.vk2ktj.ampr.org: icmp: echo request
     13:51:37.388036 gw.vk2ktj.ampr.org > albert.vk2ktj.ampr.org: icmp: echo reply
     13:51:38.383578 albert.vk2ktj.ampr.org > gw.vk2ktj.ampr.org: icmp: echo request
     13:51:38.400592 gw.vk2ktj.ampr.org > albert.vk2ktj.ampr.org: icmp: echo reply
     13:51:49.303196 albert.vk2ktj.ampr.org.1104 > gw.vk2ktj.ampr.org.telnet: S 700506986:700506986(0) win 512 <mss 1436>
     13:51:49.363933 albert.vk2ktj.ampr.org.1104 > gw.vk2ktj.ampr.org.telnet: . ack 1103372289 win 14261
     13:51:49.367328 gw.vk2ktj.ampr.org.telnet > albert.vk2ktj.ampr.org.1104: S 1103372288:1103372288(0) ack 700506987 win 2048 <mss 432>
     13:51:49.391800 albert.vk2ktj.ampr.org.1104 > gw.vk2ktj.ampr.org.telnet: . ack 134 win 14198
     13:51:49.394524 gw.vk2ktj.ampr.org.telnet > albert.vk2ktj.ampr.org.1104: P 1:134(133) ack 1 win 2048
     13:51:49.524930 albert.vk2ktj.ampr.org.1104 > gw.vk2ktj.ampr.org.telnet: P 1:28(27) ack 134 win 14335
      ..
     #

When you start tcpdump without arguments it grabs the first (lowest numbered) network device that is not the loopback device. You can specify which device to monitor with a command line argument as shown above. tcpdump then decodes each datagram transmitted or received and displays them, one line each, in a textual form. The first column is obviously the time the datagram was transmitted or received. The remainder of the line is then dependent on the type of datagram. The first two lines in the sample are what an arp request from albert.vk2ktj for gw.vk2ktj look like. The next four lines are two pings from albert.vk2ktj to gw.vk2ktj, note that tcpdump actually tells you the name of the icmp datagram transmitted or received. The greater-than (>) symbol tells you which way the datagram was transmitted, that is, from who, to who.

It points from the sender, to the receiver. The remainder of the sample trace are the establishment of a telnet connection from albert.vk2ktj to gw.vk2ktj.

The number or name at the end of each hostname tells you what socket number is being used. tcpdump looks in your /etc/services file to do this translation.

tcpdump explodes each of the fields, and so you can see the values of the window and mss parameters in some of the datagrams.

The man page documents all of the options available to you.

Note for PPP users: The version of tcpdump that is currently available does not support the PPP suite of protocols. Al Longyear has produced a set of patches to correct this, but these have not been built into a tcpdump distribution yet.

15 Some Frequently Asked Questions, with brief Answers.

Following are some questions and answers that are commonly asked.

15.1 General questions:

I have only a dialin terminal access to a machine on the net, can I use this as a network connection ?

Yes you can, take a look at TERM. TERM allows you you to run network connections over a normal terminal session. It requires some modifications to the network applications to work with it, but binaries and sources are available for the most common ones already. take a look at the TERM-

     HOWTO (http://sunsite.unc.edu/mdw/HOWTO/Term-HOWTO.html) for lots more information.

Why, when I telnet/ftp/rlogin to my machine does it take so long to answer?

You do not have your name resolver configured properly. Reread the section on /etc/resolv.conf.

I want to build my own standalone network, what addresses do I use ?

RFC1597 has specifically reserved some IP addresses for private networks. You should use these as they prevent anything nasty happening if you accidentally get connected to the Internet. The addresses reserved are:

          10.0.0.0        -   10.255.255.255
          172.16.0.0      -   172.31.255.255
          192.168.0.0     -   192.168.255.255

Note, reserved network addresses are of classes A, B and C, so you are not restricted in your network design or size. Since you won’t be connecting to the Internet it doesn’t matter if you use the same address as some other group or network, just so long as the addresses you use are unique within your network.

If sunacm.swan.ac.uk is down, how do I get the files specified ?

‘sunacm’ is mirrored on: ftp.Uni-Mainz.DE

     /pub/Linux/packages/Net2Debugged

How do I know what version of kernel/net code I am running ?

The network code and kernel now have synchronised version numbers, so try:

uname -a

or:

cat /proc/version

15.2 Error messages:

I keep getting the error ‘eth0: transmit timed out’. What does this mean?

This usually means that your Ethernet cable is unplugged, or that the setup parameters for your card (I/O address, IRQ, etc.) are not set correctly. Check the messages at boot time and make sure that your card is recognized with the correct Ethernet address. If it is, check that there is no conflict with any other hardware in your machine, eg you might have a soundblaster sharing the same IRQ or i/o control port.

I get errors ‘check Ethernet cable’ when using the network.

You probably have your Ethernet card configured incorrectly. Double check the settings in /usr/src/linux/drivers/net/CONFIG. If this checks out ok, you may in fact have a cabling problem, check the cables are plugged in securely.

15.3 Routing questions:

Why do I get the message ‘obselete route request’ when I use the route command ? You are using a version of route that is older than your kernel. You should upgrade to a newer version of route. Refer to the ”The network configuration tool suite” section of this document for information on where to obtain the tool set.

Why do I get a ‘network unreachable’ message when I try and network?

This message means that yours, or some other, machine doesn’t know how to route to the host that you are attempting to ping or connect to. If it occurs for all hosts that you try, then it is probable that you don’t have your default route set up properly, reread the ‘routing’ section.

I can ping my server/gateway, but can’t ping or connect to anyone remote.

This is probably due to a routing problem. Reread the ‘routing’ section in this document. If this looks ok, then make sure that the host you are attempting to connect to has a route to you. If you are a dialin user then this is a common cause of problems, ensure that your server is either running a routing program like gated or routed, or that it is ‘prox arping’ for you, otherwise you will be able to get datagrams to the remote host, but it won’t know how to return datagrams to you.

15.4 Using Linux with fileservers/NFS:

How do I use my existing Novell fileserver with my Linux machine ?

If you have the Novell NFS Daemon code then it is easy, just NFS mount the Novell volume that you wish to use. If you don’t, and you are really desperate to be able to do this, and you have a spare pc machine laying about, you are in luck. You can run a program called Stan’s Own Server on the spare PC. First, configure the pc as a novell workstation with maps to the directories you want to nfs mount, then run SOS, and export those drive maps. SOS is available from spdcc.com:pub/sos/sossexe.zoo

Files get corrupted when running NFS over a network.

Certain vendors (Sun primarily) shipped many machines running NFS without UDP checksums. Great on ethernet, suicide otherwise. UDP checksums can be enabled on most file servers. Linux has it enabled by default from pl13 onwards - but both ends need to have it enabled...

Why are my NFS files all read only ?

The Linux NFS server defaults to read only. RTFM the ‘exports’ and nfsd manual pages. With non Linux servers you may also need to alter /etc/exports

15.5 slip questions:

What do I do if I don’t know my slip servers address ?

dip doesn’t really need to know the address of your slip server for slip to function. The remote option was added as a convenience so that dip could automate the ifconfig and route commands for you. If you don’t know, and cannot find out the address of your slip server then Peter D. Junger Junger@samsara.law.cwru.edu has suggested that he simply used his own address wherever a dip script called for a remote address. This is a small kludge but it works ok, as the server’s address never actually appears in the slip headers anyway.

‘dip’ only works for root. How do I make it work for others?

dip needs to be setuid root to do some of the things it needs to do, such as modifying the routing table. Uri Blumenthal recommends the following:

• Create a new group called dip in your /etc/group file, and place each person who you want to allow dial out operation in it.

• Then when logged in as root, do the following:

    # chown root.dip /usr/bin/dip
    # chmod u=rx,g=x,o= /usr/bin/dip
    # chmod u+s /usr/bin/dip

Dial-In users will be restricted in what they can do by what is contained in the /etc/diphosts file.

With SLIP I can ping my server, and other hosts, but telnet or ftp don’t work. This is most likely caused by a disagreement on the use of header compression between your server and your machine. Double check that both ends either are, or are not, using compression. They must match.

How can I hang up the phone line when I’m done using SLIP?

If you use dip to dial out on the SLIP line, just ‘dip -k’ should do the trick. If not, try to kill the dip process that is running. When dip dies it should hang up the call. To give it the best chance to clean up after itself, try killing the process in the following sequence: ‘kill <pid¿’, ‘kill -hup <pid¿’, and finally, if the dip process still refuses to die, try ‘kill -9 <pid¿’. The same philosophy should be applied to all unix processes that you are attempting to kill.

I see a lot of overrun errors on my slip port, why ?

The older network tools incorrectly report number of packets compressed as the number of packets overrun. This has been corrected, and shouldn’t occur of you are running the new version kernel and tools. If it still is it probably indicates that your machine isn’t keeping up with the rate of data incoming. If you are not using 16550AFN UARTs then you should upgrade to them. 16450, or 8250 generate an interrupt for every character they receive and are therefore very reliant on the processor to be able to find time to stop what it is doing an collect the character from them to ensure none get lost. The 16500AFN has a 16 character FIFO, and they only generate interrupts when the FIFO is nearly full, or when they have had character waiting, this means that less interrupts get generated for the same amount of data, and that less time is spent servicing your serial port. If you want to use multiple serial ports you should mandatorily upgrade to 16550AFN UARTs anyway.

Can I use two slip interfaces ?

Yes. If you have, for example, three machines which you would like to interconnect, then you most certainly could use two slip interfaces on one machine and connect each of the other machines to it. Simply configure the second interface as you did the first. NOTE that the second interface will require a different IP address to the first. You may need to play with the routing a bit to get it to do what you want, but it should work.

I have a multiport i/o card, how do I use more than 4 slip ports ?

The kernel slip comes with a default of a maximum of 4 slip devices configured, this is set in the

/usr/src/linux/drivers/net/slip.h file. To increase it, say to 16, change the #define SL NRUNIT

to 16, in place of the 4 that will be there. You also need to edit /usr/src/linux/drivers/net/Space.c and add sections for sl4, sl5 etc. You can copy the existing driver definition as a template to make it easier. You will need to recompile the kernel for the change to take effect.

15.6 PPP questions.

You should refer to the PPP-HOWTO (http://sunsite.unc.edu/mdw/HOWTO/PPP-HOWTO.html) for a list

of PPP questions and answers compiled by Al Longyear.

16 Known Bugs.

The Linux networking code is still an evolving thing. It still has bugs though they are becoming less frequently reported now. The Linux Networking News (http://iifeak.swan.ac.uk/NetNews.html) is a World Wide Web page maintained by Alan Cox which contains information on the status of the NET-3 networking code. You can obtain information on what is known and what isn’t, by reading the /usr/src/linux/net/inet/README file that accompanies the kernel source, or by joining the NET channel.

17 Copyright Message.

The NET-2-HOWTO is copyright by Terry Dawson and Matt Welsh. A verbatim copy of this document may be reproduced and distributed in any medium, physical or electronic without permission of the authors. Translations are similarly permitted without express permission if such translations include a notice stating who performed the translation, and that it is a translation. Commercial redistribution is allowed and encouraged, however, the authors would like to be notified of any such distributions.

Short quotes may be used without prior consent by the authors. Derivative works and partial distributions of the NET-2-HOWTO must include either a verbatim copy of this file, or make a verbatim copy of this file available. If the latter is the case, a pointer to the verbatim copy must be stated at a clearly visible place.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we wish to retain copyright on this HOWTO document, and would like to be notified of any plans to redistribute it. Further we desire that ALL information provided in this HOWTO be disseminated.

If you have any questions relating to the conditions of this copyright, please contact Matt Welsh, the Linux HOWTO coordinator, at: mdw@sunsite.unc.edu

18 Miscellaneous, and Acknowledgements.

This HOWTO has been completely rewritten using the new smgl tools that Matt Welsh put together. The tools seem to work just fine, and they are pretty simple to use. There are so many people who have contributed comments and suggestions for this update that I have forgotten who you are. Thanks.

Please, if you have any comments or suggestions then mail them to me. I’m fairly busy these days, so I might not get back to you straight away, but I will certainly consider any suggestion you have.

The Linux networking code has come a long way, and it hasn’t been an easy trip, but the developers, all of them, have done an excellent job in getting together something that is functional, versatile, flexible, and free for us to use. We all owe them a great debt of thanks. Linus, Ross, Fred, Alan, the Alpha/Beta testers, the tools developers, and those offering moral support have all contributed to the code as it is today.

For those that have an itch they want to scratch, happy hacking, here it is.

regards Terry Dawson, vk2ktj.
<terryd@extro.ucc.su.oz.au>, or <terry@orac.dn.telecom.com.au>